### v1

# Frequency-Explicit Shape Holomorphy in Uncertainty Quantification for Acoustic Scattering

Abstract
We consider frequency-domain acoustic scattering at a homogeneous star-shaped penetrable obstacle, whose shape is uncertain and modelled via a radial spectral parameterization with random coefficients. Using recent results on the stability of Helmholtz transmission problems with piecewise constant coefficients from [[[[A. Moiola and E. A. Spence, Acoustic transmission problems: wavenumber-explicit bounds and resonance-free regions, Mathematical Models and Methods in Applied Sciences, 29 (2019), pp. 317–354]]]] we obtain frequency-explicit statements on the holomorphic dependence of the scattered field and the far-field pattern on the stochastic shape parameters. This paves the way for applying general results on the efficient construction of high-dimensional surrogate models. We also take into account the effect of domain truncation by means of perfectly matched layers (PML). In addition, spatial regularity estimates which are explicit in terms of the wavenumber $k$ permit us to quantify the impact of finite-element Galerkin discretization using high-order Lagrangian finite-element spaces.
1 Introduction
1.1 Scattering Transmission Problem
We consider frequency-domain acoustic scattering by a homogeneous penetrable scatterer occupying an open, bounded set $D\subset\mathbb{R}^{d}$, $d=2,3$, with Lipschitz boundary $\partial D$, which is embedded in a homogeneous background medium occupying $\mathbb{R}^{d}\backslash\overline{D}$. This can be modeled mathematically by a transmission problem for the Helmholtz equation with the Sommerfeld radiation conditions at infinity [^1] Vectors in Euclidean space are denoted by bold roman symbols, ${\boldsymbol{x}}\in\mathbb{R}^{3}$, and $|{\boldsymbol{x}}|$ stands for its Euclidean norm.:
$\big{(}-k^{-2}\Delta-n({\boldsymbol{x}})\big{)}u({\boldsymbol{x}})$ $=0$ $\text{in}\;\mathbb{R}^{d},$ (1.1a)
[…]
Here, $k>0$ is the wavenumber, which is proportional to the angular frequency, and $n=n({\boldsymbol{x}})$ is a spatially varying, but piecewise constant, index of refraction, for which we assume
$n({\boldsymbol{x}})=\begin{cases}n_{i}>0&\text{for }{\boldsymbol{x}}\in D,\\ 1&\text{for }{\boldsymbol{x}}\in\mathbb{R}^{d}\setminus\overline{D}.\end{cases}$ (1.2)
In the acoustic modeling context, the solution $u$ of (1.1) gives the complex amplitude of the sound pressure of the so-called scattered wave, see [22, Sect 2.1] for more detail.
Excitation is provided by an incident wave $u^{\mathrm{inc}}$ that satisfies $-\Delta u^{\mathrm{inc}}-k^{2}u^{\mathrm{inc}}=0$ in $\mathbb{R}^{d}$, the prime example being a plane wave $u^{\mathrm{inc}}({\boldsymbol{x}})=\exp({\rm i}k{\boldsymbol{d}}\cdot{\boldsymbol{x}})$, propagating in direction ${\boldsymbol{d}}\in\mathbb{R}^{d}$, $|{\boldsymbol{d}}|=1$. Existence and uniqueness of a solution $u\in H^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$ of (1.1) is well-known, see, e.g., [53, Lemma 2.2].
In this article we focus on the model problem (1.1)–(1.2) for the sake of simplicity. Our analysis could also be extended to settings with more general piecewise constant coefficient functions in the zero-order and second-order terms of the Helmholtz equation (1.1a), which would not require fundamentally new ideas.
1.2 Uncertainty Quantification (UQ) via Polynomial Surrogate Models
The shape of the scatterer $D$ is assumed to be uncertain, which we take into account by the customary approach to deterministic UQ through stochastic parameterization: the uncertainty in $D$ is captured by introducing a dependence of $D$ on (possibly infinitely many) real-valued random variables with known distributions. The key idea is to relegate those to mere real parameters. Thus we arrive at a family of transmission problems depending on infinitely many parameters. The solution $u$ of the transmission problem and any derived quantity of interest will then become functions of the parameters, alike.
Next we build sparse polynomial surrogate models of those functions on the parameter domain. This can be done accurately and efficiently using suitable spaces of multi-variate polynomials in the parameters, provided that $u$ and the quantities of interest are analytic/holomorphic [^2] We use “analytic” and “holomorphic” as synonyms, cf. [55, Definition 5.1]. functions of the parameters with a sufficiently large domain of analyticity in the complexified parameter space.
1.3 Simplest Case: Size Uncertainty Quantification
We first consider the case that only the size of scatterer $D$ and not its shape is random, and model it by setting
$D=D(\omega):=(1+\tfrac{1}{2}Y(\omega))D_{0}\;,\quad\omega\in\Omega,$ (1.3)
where
  • $D_{0}\subset\mathbb{R}^{d}$ is a bounded Lipschitz domain (the “nominal scatterer”), and
  • $Y:\Omega\to[-1,1]$ is a random variable on some probability space $\Omega$, whose details are not important at this stage.
We follow the policy outlined in the previous section and replace $Y(\omega)$ with a single real parameter $y\in[{-1,1}]$. This yields the parametric domain model
$D(y):=(1+\tfrac{1}{2}y)D_{0}\;,\quad-1\leq y\leq 1\;.$ (1.4)
Simple scaling arguments show that, if $u=u(y)\in H^{1}_{\rm loc}(\mathbb{R}^{d})$ solves the transmission problem (1.1), then $\widehat{u}=\widehat{u}(y)\in H^{1}_{\rm loc}(\mathbb{R}^{d})$ defined as
$\widehat{u}(y;\widehat{{\boldsymbol{x}}}):=u((1+\tfrac{1}{2}y)\widehat{{\boldsymbol{x}}})\;,\quad\widehat{{\boldsymbol{x}}}\in\mathbb{R}^{d}\;,$ (1.5)
solves (1.1) with (i) $D(y)$ replaced with $D_{0}$, (ii) a modified wave number $\widehat{k}=\widehat{k}(y):=(1+\tfrac{1}{2}y)k$, and (iii) the exciting field $\widehat{u}_{\mathrm{inc}}(\widehat{{\boldsymbol{x}}}):=u_{\mathrm{inc}}((1+\tfrac{1}{2}y)\widehat{{\boldsymbol{x}}})$. We realize that, up to an affine transformation, the dependence of $\widehat{u}$ on the scaling parameter $y$ is the same as its dependence on the wave number $k$.
The dependence of the solution of the Helmholtz transmission problem on the wave number $k$ is a classic topic of study in scattering theory. One key result is when $n_{i}>1$ and $D_{0}$ is smooth with strictly-positive curvature the norm of the solution operator grows superalgebraically through an increasing sequence of wavenumbers [57]  [^3] Strictly speaking, [57] proves the existence of a sequence of resonances exponentially close to the real axis, but then the “resonances to quasimodes” result of [62] implies super-algebraic growth through a sequence of real $k_{j}$s.; i.e., in the scaled setting above, there exist sequences of real wavenumbers $(\widehat{k}_{j})_{j\in\mathbb{N}}$, $\widehat{k}_{j}\to\infty$ for $j\to\infty$, such that, given $R,N>0$, there exists $C_{N}>0$ such that [^4] We write $B_{R}$ for the open ball around 00 in $\mathbb{R}^{d}$.
$\left\|{\widehat{u}}\right\|_{H^{1}(B_{R})}\geq C_{N}(\widehat{k}_{j})^{N}\quad\text{ for all }j\in\mathbb{N}.$ (1.6)
At least when $D_{0}$ is a ball, this growth is exponential, i.e., there exist $C,\gamma>0$ such that
$\left\|{\widehat{u}}\right\|_{H^{1}(B_{R})}\geq C\exp\big{(}\gamma\widehat{k}_{j}\big{)}\quad\text{ for all }j\in\mathbb{N};$ (1.7)
see [6, 7, 1]. To explain this growth, recall that resonances of the Helmholtz transmission problem are poles of the meromorphic extension $\widehat{k}\in\mathbb{C}\mapsto\widehat{u}(\widehat{k})$. When $n_{i}>1$ and $D_{0}$ is smooth with strictly-positive curvature there are resonances superalgebraically close to the real axis [57, Theorem 1.1], corresponding physically to “whispering-gallery modes” created by total internal reflection of rays hitting $\partial D_{0}$ from $D_{0}$. The $\widehat{k}_{j}$, often called quasi-resonances, can then be thought of as the real parts of these resonances close to the real axis; see Figure 1.1 (a) for a numerical illustration. We highlight also that the density of the near-real-axis resonances, and hence also the $\widehat{k_{j}}$, increases as $j$ increases [9, Theorem 1.3].
(a) $n_{i}=3$
(b) $n_{i}=\frac{1}{3}$
Figure 1.1: Dependence of norms ${\left\|{u}\right\|_{L^{2}(B_{2})}}$ and ${\left\|{u}\right\|_{H^{1}(B_{2})}}$ of the solution $u$ of the transmission problem (1.1) on $k$, when $d=2$, $D$ is the unit disk, and $u_{\mathrm{inc}}({\boldsymbol{x}})=\exp(\rm ikx_{1})$. For $n_{i}=3$ quasi-resonances manifest themselves as spikes of the graph $k\mapsto\left\|{u}\right\|_{H^{1}(B_{2})}$, for $n_{i}=\frac{1}{3}$ such spikes are conspicuously absent. The norms were computed by means of a Fourier spectral method with a number of modes large enough to render any discretization error negligible. MATLAB codes in https://github.com/hiptmair/ScatteringQuasiResonances.
From the relationship $\widehat{k}(y)=(1+\frac{1}{2}y)k$ we conclude that the mapping $y\in[{-1,1}]\mapsto\left\|{\widehat{u}}\right\|_{H^{1}(D_{0})}$ may feature spikes, which will become more numerous, steeper, and higher with increasing $k$. Therefore, although the function $y\mapsto\widehat{u}(y)$ is analytic, already for moderate $k$ its accurate polynomial approximation may require very high degrees. Putting it bluntly, polynomial surrogate modeling is doomed in the presence of quasi-resonances and at this point no viable alternative is available, which forces us to avoid scattering problems beset with quasi-resonances.
Remark 1.1.
A promising approach to obtain efficient surrogate models for the map $\widehat{k}\mapsto\widehat{u}(k)$ may rely on rational functions, see [3]. So far this is confined to a single parameter and extension to many parameters as required by general shape UQ methods is not clear yet.
1.4 Scattering problems without quasi-resonances
Whereas the Helmholtz transmission problem with $n_{i}>1$ can suffer from quasi-resonances, the problem with $n_{i}<1$ does not. Indeed, the condition $n_{i}<1$ rules out total internal reflection of rays hitting $\partial D_{0}$ from $D_{0}$. Furthermore [53] showed that when $n_{i}<1$ and $D$ is star-shaped Lipschitz, given $k_{0},R>0$, for all $k\geq k_{0}$, $k_{0}>0$ sufficiently large, the outgoing solution of
${(k^{-2}\Delta+n)u=f}\quad\text{ with }\quad f\in L^{2}_{\rm comp}(B_{R})$ (1.8)
satisfies
$\left\|{u}\right\|_{H^{1}_{k}(B_{R})}\leq Ck\left\|{f}\right\|_{L^{2}({B_{R}})}\;,$ (1.9)
where $\left\|{u}\right\|^{2}_{H^{1}_{k}(B_{R})}:=k^{-2}\left\|{\nabla u}\right\|^{2}_{L^{2}(B_{R})}+\left\|{u}\right\|^{2}_{L^{2}(B_{R})}$ and $C$ depends only on $n_{i}$, $d$, and $R$; i.e.,
the $k$-explicit bound (1.9) holds uniformly on the set of star-shaped scatterers.
This observation motivated the present paper, because star-shapedness is a simple geometric property, whose persistence under random shape perturbation seems rather natural. The bound (1.9) implies that the poles of the meromorphic function $k\in\mathbb{C}\mapsto u(k)$ are located below $\left\{z\in\mathbb{C}:\,\operatorname{Im}z=-\nu\right\}$ for some $\nu>0$, i.e., there is a resonance-free strip beneath the real axis, which means that no quasi-resonances are present, see Figure 1.1 (b). This resonance-free-strip result was proved for smooth $D$ with strictly-positive curvature in [8], and more-detailed information about the location of the resonances in this case is given in [30].
We now explore the implications of the results from [53] in the simple size-UQ setting of Section 1.3. If $n_{i}\leq 1$ and $D_{0}$ is star-shaped with respect to 00, then $y\in\mathbb{C}\mapsto\widehat{u}(y)$ is meromorphic on $\left\{z\in\mathbb{C}:\,\operatorname{Re}z>{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}{-2}}\right\}$ 2 }. However, since $\widehat{k}(y)=(1+\frac{1}{2}y)k$, the distance of the poles of $y\mapsto\widehat{u}(y)$ to the real axis shrinks like $O(k^{-1})$ as $k\to\infty$. This means that the extent of the domain of analyticity of the function $y\in\mathbb{C}\mapsto\widehat{u}(y)$ in the direction of the imaginary axis will decrease like $O(k^{-1})$ for $k\to\infty$. Therefore, even in non-resonant situations, $k$-uniformly accurate polynomial surrogate models for $y\mapsto\widehat{u}(y)$ will require increasingly larger degree for growing $k$. This growth with respect to $k$ can be avoided only if the size of admissible shape perturbations (controlled by the range of the parameter $y$ in the size-UQ model) decreases as $O(k^{-1})$ for $k\to\infty$. In other words,
shape perturbations must be confined to the scale of the wavelength (or smaller) to permit provably $k$-robust polynomial surrogate modelling.
Remark 1.2.
We highlight that this condition of $O(k^{-1})$ perturbations is encountered in the UQ of the Helmholtz equation with variable coefficients, i.e.,
$k^{-2}\nabla_{\boldsymbol{x}}\cdot\big{(}A({\boldsymbol{x}},{\boldsymbol{y}})\nabla_{\boldsymbol{x}}u({\boldsymbol{x}},{\boldsymbol{y}})\big{)}+n({\boldsymbol{x}},{\boldsymbol{y}})u({\boldsymbol{x}},{\boldsymbol{y}})={f}({\boldsymbol{x}})$
with $u$ satisfying the Sommerfeld radiation condition and $A({\boldsymbol{x}},{\boldsymbol{y}})$ and $n({\boldsymbol{x}},{\boldsymbol{y}})$ perturbations of some $A_{0}({\boldsymbol{x}})$ and $n_{0}({\boldsymbol{x}})$, respectively. When the problem with $A_{0}$ and $n_{0}$ is nontrapping, given $k_{0}>0$ there exists $C>0$ such that for all $k\geq k_{0}$ the map ${\boldsymbol{y}}\mapsto u({\boldsymbol{y}})$ is analytic for all ${\boldsymbol{y}}$ such that [61, Theorem 1.1]
$k\max\big{\{}\|A-A_{0}\|_{L^{\infty}},\|n-n_{0}\|_{L^{\infty}}\big{\}}\leq C\;,$ (1.10)
with this bound sharp through an unbounded sequence of wavenumbers $k$ [61, Theorem 1.4]. In [34, Theorem 4.2], the condition (1.10) is also shown to be a sufficient condition for analyticity (in the form of the relevant bounds on derivatives of $u$ with respect to ${\boldsymbol{y}}$) in the case when $A\equiv I$, the Sommerfeld radiation condition is approximated by an impedance boundary condition, and for particular classes of nontrapping $n_{0}$ and perturbations $n-n_{0}$.
1.5 Layout of the paper
The foundations for the current work were laid in [41], which dealt with shape UQ for the Helmholtz transmission problem (1.1), (1.2) in a non-$k$-explicit way under the assumption that the wave number was smaller than any possible quasi-resonance. Here we reuse a tool of [41], the parameterization of the shape of $D$ by a radial displacement function $r\in C^{1}(\mathbb{S}^{d-1})$ defined on the unit sphere $\mathbb{S}^{d-1}$ in $\mathbb{R}^{d}$:
[…]
If $\left\|{r}\right\|_{C^{0}(\mathbb{S}^{d-1})}<1$, then any $D(r)$ is star-shaped with respect to 00 and if, in addition, $n_{i}\leq 1$, then the requirements of [53] for non-resonant scattering are satisfied immediately.
The shape parameterization (1.11) is examined in Section 2, where we extend it to a mapping of $D(r)$ to the unit sphere $B_{1}$ to convert (1.1) into a variational problem with $r$-dependent coefficients. This is the gist of the popular domain mapping method for shape UQ, first proposed in [64] and then used by many other authors, e.g. in [38, 10, 16]. Yet, the latter works rely on volume diffeomorphisms for parameterization of shapes, which do not offer a natural avenue to the preservation of star-shapedness.
In Section 3 we review the results of [53], apply them to the concrete transmission problem (1.1)-(1.2) and state more precisely the estimate (1.9) and generalizations of it, providing completely $k$-explicit expressions for the constants. Then, in Section 4, we admit $\mathbb{C}$-valued radial displacement functions $r$ in (1.11); thanks to the domain mapping approach the resulting transmission problem remains well defined and a perturbation argument yields $k$-explicit stability estimates. These estimates give detailed $k$-explicit information about shape holomorphy, more precisely, about the domain of analyticity of $r\mapsto\widehat{u}(r)$, $\widehat{u}(r)$ the solution of (1.1), with $D=D(r)$ given by (1.11), pulled back to the unit sphere. Such “shape holomorphy” results have been established in recent years for a host of parametric PDEs models [10, 2, 39, 40, 19, 45]. In these works, the size of the holomorphy domain of parametric solutions and operators was not made explicit in terms of problem parameters such as wavenumber, or (in [19]) the Reynolds number.
Next, in Section 5 we investigate the impact of a high-order finite-element discretization of the transmission problem. We give $k$-explicit estimates of how the finite-element discretization error affects the estimates of statistical moments.
In Section 6 we present the so-called far-field pattern as a representative of relevant output functionals that depend on $\widehat{u}(r)$ and inherit its holomorphy.
In Section 7 we introduce a second-level affine parameterization of $r$ through a sequence of uniformly-distributed random variables. Then, in Section 8, we appeal to the theory of [16, 65, 66], and leverage the available information on shape holomorphy to predict the rate of convergence of Smolyak-type, high-dimensional quadrature when applied to extract the statistical mean of the far-field-pattern random field.
List of notations
$d$$\hat{=}$ spatial dimension $\in\{2,3\}$
$k$$\hat{=}$ positive wavenumber
$r=r(\widehat{{\boldsymbol{s}}})$$\hat{=}$ radial displacement function $\in C^{1}(\mathbb{S}^{d-1})$, see (1.11)
$u=u(r;\widehat{{\boldsymbol{x}}})$$\hat{=}$ solution of scattering problem, see (2.7)
$\boldsymbol{\Phi}=\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$$\hat{=}$ transformation from reference domain, see (2.3)
$\chi=\chi(\rho)$$\hat{=}$ cut-off function defined in (2.4)
$\widehat{u}=\widehat{u}(r;\widehat{{\boldsymbol{x}}})$$\hat{=}$ solution of transformed scattering problem (2.8)
$\widehat{{\mathbf{A}}}=\widehat{{\mathbf{A}}}(r;\widehat{{\boldsymbol{x}}})$$\hat{=}$ diffusion coefficient tensor of transformed variational problem, see (2.9a)
$\widehat{n}=\widehat{n}(r;\widehat{{\boldsymbol{x}}})$$\hat{=}$ refractive index of transformed variational problem, see (2.9b)
$\left\|{\cdot}\right\|_{H^{1}_{k}(B_{2})}$, $\left\|{\cdot}\right\|_{H^{2}_{k}(B_{2})}$$\hat{=}$ $k$-scaled Sobolev norms, see (3.1)
$\mathfrak{r}=\mathfrak{r}(\widehat{{\boldsymbol{s}}})$$\hat{=}$ complex-valued radial displacement function $\in C^{1}(\mathbb{S}^{d-1},\mathbb{C})$
$\mathfrak{R}$$\hat{=}$ set of admissible complex-valued radial displacement functions, see (4.1)
$\widehat{a}=\widehat{a}(\mathfrak{r};\widehat{u},\widehat{v})$$\hat{=}$ bilinear form of transformed variational problem, see (4.3)
${\cal A}={\cal A}(k;C_{\operatorname{Re}},C_{\operatorname{Im}})$$\hat{=}$ domain of analyticity of $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})$, see (4.7)
$\widehat{u}_{\rm PML}=\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$$\hat{=}$ solution of PML-truncated transformed scattering problem
$\Omega_{\mathrm{tr}}$$\hat{=}$ PML-truncated computational domain
$\widetilde{\sigma}=\widetilde{\sigma}(\widehat{(}\rho))$$\hat{=}$ PML control function, see (5.3)
$p\in\mathbb{N}$$\hat{=}$ polynomial degree/approximation order of finite element space, see Assumption 5.7
$V_{h}$$\hat{=}$ finite element space $\subset H^{1}(\Omega_{\mathrm{tr}})$
$\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}\in V_{h}$$\hat{=}$ Galerkin FE solution with PML truncation
${\mathcal{A}}_{p}={\mathcal{A}}_{p}(k;C_{\operatorname{Re}},C_{\operatorname{Im}})$$\hat{=}$ set of admissible complex-valued radial displacement function for FEM, see (5.10)
$\widehat{u}^{\rm scat}_{\infty}=\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};{\boldsymbol{q}})$$\hat{=}$ far-field pattern, see (6.8)
$\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$$\hat{=}$ approximate far-field pattern based on FEM and PML, see (6.11)
$r_{j}=r_{j}({\boldsymbol{x}})$$\hat{=}$ radial expansion functions, see (7.2)
${\cal P}$$\hat{=}$ set of shape parameter sequences, see (7.3)
${\cal H}={\cal H}(C_{\operatorname{Re}},C_{\operatorname{Im}},k)$$\hat{=}$ domain of analyticity of ${\mathbf{z}}\mapsto\breve{u}({{\mathbf{z}}};\cdot)$, see Corollary 7.4.
2 Variational Formulation of the Transmission Problem with Parametric Interface
The parameterization (1.11), apart from nicely fitting the assumptions of the theory of [53], also paves the way for a detailed and explicit analysis, on which we now embark.
2.1 A parametric family of diffeomorphisms
Recall the parameterization of the shape of the scatterer $D$ by means of a continuously differentiable radial displacement function $r\in C^{1}(\mathbb{S}^{d-1},\mathbb{R})$ defined on the $d$-dimensional unit sphere $\mathbb{S}^{d-1}\subset\mathbb{R}^{d}$
[…]
It goes without saying that $r$ has to be bounded from below at least to render (1.11) meaningful. For our analysis we confine $r$ to a more narrow range of values:
Assumption 2.1 (Bounds for radial displacement function).
We admit only radial displacement functions belonging to the set
:=${\cal R}:=\left\{r\in C^{1}(\mathbb{S}^{d-1},\mathbb{R}),\,\left\|{r}\right\|_{C^{0}(\mathbb{S}^{d-1})}\leq\frac{1}{3}\right\},$ (2.1)
that is $|r({\widehat{{\boldsymbol{s}}}})|\leq\frac{1}{3}$ for all $\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1}$.
Note that if $r\in C^{m}(\mathbb{S}^{d-1},\mathbb{R})$, then $D(r)$ is of class $C^{m}$. In particular, $D(r)$ is a bounded Lipschitz domain and the interface $\Gamma(r):=\partial D(r)$ is contained in a spherical shell:
[…]
As in [41, Section 3], we shift parameter dependence from the domain $D(r)$ to the coefficients of a transformed transmission problem, for which the scatterer occupies the unit ball $\widehat{D}:=\{{\boldsymbol{x}}\in\mathbb{R}^{d}:\,|{\boldsymbol{x}}|<1\}$. Taking the cue from [41, Equation (3.2)] or [35, Section 3.2], the transformation is effected by the parameter-dependent mapping $\boldsymbol{\Phi}(r):\mathbb{R}^{2}\to\mathbb{R}^{2}$, $r\in{\cal R}$, given by
$\boxed{\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})\mid_{\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in\mathbb{R}^{d}}=\widehat{{\boldsymbol{s}}}\bigl{(}\widehat{\rho}+\chi(\widehat{\rho})r(\widehat{{\boldsymbol{s}}})\bigr{)}\;,\quad\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1}\,,\;\widehat{\rho}\geq 0\;,\quad\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in\mathbb{R}^{d}}\;,$ (2.3)
with a cut-off function [^5] We can choose $\chi(\rho)=f(g(\frac{\rho-1}{1-\lambda}))$ with $f(\xi):=\exp\left(1-\frac{1}{1-\xi^{2}}\right)$ for $|\xi|<1$, $f(\xi):=0$ for $|\xi|\geq 1$, $g(\zeta):=\arctan(\frac{3}{2}\zeta)/\arctan(\frac{3}{2})$, and $\lambda=\frac{1}{40}$. This function is displayed in the plot beside. $\chi\in C^{\infty}([{0,\infty}[)$, satisfying for some $0<\lambda\ll 1$
(i) $0\leq\chi(\rho)$ $\leq 1\quad\forall\rho\geq 0,$ (2.4a)
(ii) $\chi(\rho)$ $=0\quad\text{for}\;\rho\leq\lambda\;\text{or}\;\rho\geq 2-\lambda,$ (2.4b)
(iii) $|\chi^{\prime}(\rho)|$ $\leq{2-\lambda}\quad\forall\rho\geq 0,$ (2.4c)
(iv) $\chi(1)$ $=1.$ (2.4d)
Property (2.4c) combined with (2.2) ensures that the radial dilation effected by $\boldsymbol{\Phi}(r;\cdot)$, $\rho\mapsto\rho+\chi(\rho)r(\widehat{\widehat{{\boldsymbol{s}}}})$, has a derivative bounded below by $\frac{1}{3}$ and, hence, is strictly monotone. We therefore have the following lemma.
Lemma 2.2 (Properties of mappings $\boldsymbol{\Phi}$).
Under Assumption 2.1 and (2.4) the mappings $\boldsymbol{\Phi}(r;\cdot)$, $r\in{\cal R}$, as defined in (2.3) are diffeomorphisms of class $C^{m-1,1}$, map the unit sphere $\partial\widehat{D}=\mathbb{S}^{d-1}$ onto the interface $\Gamma(r)$, and agree with the identity in the exterior of the spherical shell $\{\widehat{{\boldsymbol{x}}}\in\mathbb{R}^{d}:\,\lambda<|\widehat{{\boldsymbol{x}}}|<2-\lambda\}$.
2.2 Mapped variational formulation of
Recall that for all $r\in{\cal R}$ the domain $D(r)$ is contained in the ball $B_{2}$; therefore the transmission problem (1.1) can be written as a variational problem on $B_{2}$ involving the Dirichlet-to-Neumann map on $\partial B_{2}$. Given $g\in H^{1/2}(\partial B_{2})$, let $v$ be the solution of
$(-k^{-2}\Delta-1)v=0\quad\text{ in }\mathbb{R}^{d}\setminus\overline{B_{2}}\quad\text{ and }\quad v=g\text{ on }\Gamma_{2}:=\partial B_{2}$ (2.5)
satisfying the Sommerfeld radiation condition
$|{\boldsymbol{x}}|^{\frac{d-1}{2}}\left(k^{-1}\frac{\partial}{\partial|{\boldsymbol{x}}|}-\rm i\right)v({\boldsymbol{x}})\to 0\quad\text{as}\;|{\boldsymbol{x}}|\to\infty,\text{uniformly in ${\boldsymbol{x}}/|{\boldsymbol{x}}|$}.$ (2.6)
Then, define the map $\mathsf{DtN}:H^{1/2}(\partial B_{2})\to H^{-1/2}(\partial B_{2})$ by
$(\mathsf{DtN}g)({\boldsymbol{x}}):=k^{-1}\nabla v({\boldsymbol{x}})\cdot{\boldsymbol{n}}({\boldsymbol{x}})\;,\quad{\boldsymbol{x}}\in\partial B_{2},$
where on $\partial B_{2}$ we have ${\boldsymbol{n}}({\boldsymbol{x}}):=R^{-1}{\boldsymbol{x}}$, i.e., ${\boldsymbol{n}}({\boldsymbol{x}})=\frac{{\boldsymbol{x}}}{|{{\boldsymbol{x}}}|}$ is the outward-pointing unit normal vector to $B_{2}$.
Through the $r$-dependent domain $D=D(r)$, the weak solution $u\in H^{1}(B_{2})$ of the transmission problem obviously depends on the radial displacement function $r\in{\cal R}$, with ${\cal R}$ defined in Assumption 2.1; we therefore write $u=u(r)$, regarding $u$ as a mapping ${\cal R}\to H^{1}(B_{2})$, which turns out to be convenient in §7. Recall that, by Assumption 2.1, $D=D(r)$ is contained in the ball $B_{2}$; therefore the variational formulation of (1.1) is:
Given $L\in(H^{1}(B_{2}))^{*}$, $r\in{\cal R}$, and piecewise-constant refractive index $n=n(r,\cdot)$ defined by (1.2) and (1.11), find $u{(r;\cdot)}\in H^{1}(B_{2})$ such that, for all $v\in H^{1}(B_{2})$, $\int_{B_{2}}\Big{(}k^{-2}\boldsymbol{\nabla}u({r};{\boldsymbol{x}})\cdot\overline{\boldsymbol{\nabla}v}({\boldsymbol{x}})-n(r;{\boldsymbol{x}})u({r};{\boldsymbol{x}})\overline{v}({\boldsymbol{x}})\Big{)}{\rm d}{\boldsymbol{x}}-\int\nolimits_{\partial B_{2}}k^{-1}\mathsf{DtN}(\left.\mathstrut u({r;\cdot})\right|_{\partial B_{2}})\overline{v}\,\mathrm{d}S=L(v).$ (2.7)
If
$L(v):=k^{-1}\int\nolimits_{\partial B_{2}}\bigl{(}k^{-1}\boldsymbol{\nabla}u^{\mathrm{inc}}\cdot{\boldsymbol{n}}-\mathsf{DtN}(\left.\mathstrut u^{\mathrm{inc}}\right|_{\partial B_{2}})\bigr{)}\,\overline{v}\,\mathrm{d}S,$
then the solution of (2.7) is the restriction to $B_{2}$ of the solution of (1.1).
We repeat that the variational problem (2.7) depends on $r\in{\cal R}$ via $D=D(r)$ and (1.2). This dependence becomes more apparent in the transformed variational problem formed by pulling back (2.7) to the nominal setting with interface $\partial\widehat{D}=\mathbb{S}^{d-1}$ through the diffeomorphism $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$. As derived in [41, Section 4.1], this transformed variational problem reads:
Find $\widehat{u}\in H^{1}(B_{2})$ such that $\int\nolimits_{B_{2}}\Big{(}k^{-2}\widehat{{\mathbf{A}}}(r;\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}(r;\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}}(\widehat{{\boldsymbol{x}}})-\widehat{n}(r;\widehat{{\boldsymbol{x}}})\widehat{u}(r;\widehat{{\boldsymbol{x}}})\overline{\widehat{v}}(\widehat{{\boldsymbol{x}}})\Big{)}{\rm d}\widehat{{\boldsymbol{x}}}-\\ \int\nolimits_{\partial B_{2}}k^{-1}\mathsf{DtN}(\left.\mathstrut\widehat{u}(r;\cdot)\right|_{\partial B_{2}})\overline{\widehat{v}}\,\mathrm{d}S=\widehat{L}(\widehat{v})\quad\forall\widehat{v}\in H^{1}(B_{2})\;,$ (2.8)
where $\widehat{{\mathbf{A}}}(r;\widehat{{\boldsymbol{x}}})$ $:=\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})^{-1}\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})^{-\top}\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}}),$ (2.9a)
$\widehat{n}(r;\widehat{{\boldsymbol{x}}})$ $:=\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})n_{0}(\widehat{{\boldsymbol{x}}})\;,\quad n_{0}(\widehat{{\boldsymbol{x}}}):=\begin{cases}n_{i}&\text{for}\;|\widehat{{\boldsymbol{x}}}|<1,\\ 1&\text{for}\;|\widehat{{\boldsymbol{x}}}|\geq 1,\end{cases}$ (2.9b)
$\widehat{L}(\widehat{v})$ $:=L(\boldsymbol{\Phi}(r)^{*}v)\;,\,\,\text{ where }\,\,(\boldsymbol{\Phi}(r)^{*}v)(\widehat{{\boldsymbol{x}}}):=v(\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})).$ (2.9c)
We recall that, by (2.3), $\boldsymbol{\Phi}(r,\widehat{{\boldsymbol{x}}}):=\widehat{{\boldsymbol{s}}}(\widehat{\rho}+\chi(\widehat{\rho})r(\widehat{{\boldsymbol{s}}}))$, $\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1}$, $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}$.
To understand how the coefficients $\widehat{{\mathbf{A}}}$ and $\widehat{n}$ depend on the function $r\in{\cal R}$, we establish their precise expressions. The following elementary lemma shows that the Jacobian $\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r,\widehat{{\boldsymbol{x}}})$ of $\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ has triangular structure with only off-diagonal depending on derivatives of $r(\widehat{{\boldsymbol{s}}})$.
Lemma 2.3 (Jacobian of transformation mapping).
Given $r\in{\cal R}$, in polar/spherical coordinates the Jacobian of the mapping $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ from (2.3) at $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in B_{2}$, $0\leq\widehat{\rho}\leq R$, $\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1}$, is
$\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})=\begin{bmatrix}1+\chi^{\prime}(\widehat{\rho})r(\widehat{{\boldsymbol{s}}})&\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}\dfrac{dr}{d\widehat{{\boldsymbol{s}}}}(\widehat{{\boldsymbol{s}}})\\ \mathbf{0}&\left(1+\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}r(\widehat{{\boldsymbol{s}}})\right)\,{\mathbf{I}}_{d-1}\end{bmatrix}\in\mathbb{R}^{d\times d}.$ (2.10)
Proof.
For fixed $r\in C^{1}(\mathbb{S}^{d-1},\mathbb{R}^{+})$ the mapping $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}}):=\widehat{{\boldsymbol{s}}}\bigl{(}\widehat{\rho}+\chi(\widehat{\rho})r(\widehat{{\boldsymbol{s}}})\bigr{)}$, $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in B_{2}\subset\mathbb{R}^{d}$, from (2.3) is of the form
$B_{2}\to\mathbb{R}^{d}\quad,\quad{\boldsymbol{x}}\mapsto\boldsymbol{\Psi}({\boldsymbol{x}}):=\psi\bigl{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\bigr{)}\,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|},$ (2.11)
with a continuously differentiable function $\psi:(0,2]\times\mathbb{S}^{d-1}\to\mathbb{R}$ with $\psi(r,\widehat{{\boldsymbol{s}}})=r$ for small $r$. Hence it suffices to compute the derivative of $\boldsymbol{\Psi}$ for ${\boldsymbol{x}}\in B_{2}\setminus\{0\}$.
Let ${\boldsymbol{h}}\in\mathbb{R}^{d}$ be a sufficiently-small perturbation vector. By Taylor expansion, as ${\boldsymbol{h}}\to 0$,
$|{\boldsymbol{x}}+{\boldsymbol{h}}|$ $=|{\boldsymbol{x}}|+\frac{{\boldsymbol{x}}^{\top}{\boldsymbol{h}}}{|{\boldsymbol{x}}|}+O(|{\boldsymbol{h}}|^{2}),$ (2.12)
$\frac{{\boldsymbol{x}}+{\boldsymbol{h}}}{|{\boldsymbol{x}}+{\boldsymbol{h}}|}$ $=\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}+\left(\frac{{\mathbf{I}}}{|{\boldsymbol{x}}|}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{3}}\right){\boldsymbol{h}}+O(|{\boldsymbol{h}}|^{2}).$ (2.13)
We write $\partial_{1}\psi$ and $\partial_{2}\psi$ for the partial derivatives of $\psi$; note that $\partial_{1}\psi$ is a scalar, and $\partial_{2}\psi$ is a $1\times d$ matrix (i.e., a row vector). Using (2.12) and (2.13) followed by Taylor expansion gives us
[…]
=$2\displaystyle=\boldsymbol{\Psi}({\boldsymbol{x}})+\begin{aligned} &\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\,\partial_{1}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\frac{{\boldsymbol{x}}^{\top}{\boldsymbol{h}}}{|{\boldsymbol{x}}|}+\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\,\partial_{2}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\frac{1}{|{\boldsymbol{x}}|}\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{2}}}}\right){\boldsymbol{h}}\\ &+\frac{1}{|{\boldsymbol{x}}|}\psi\Big{(}|{\boldsymbol{x}}|,\frac{{\boldsymbol{x}}}{|{\boldsymbol{x}}|}\Big{)}\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{2}}\right){\boldsymbol{h}}+O(|{\boldsymbol{h}}|^{2}).\end{aligned}$
Let $\mathsf{P}_{r}({\boldsymbol{x}})$ and $\mathsf{P}_{\varphi}({\boldsymbol{x}})$ denote the orthogonal projections onto ${\boldsymbol{x}}\right>$ and ${\boldsymbol{x}}\right>^{\perp}$, i.e.,
$\mathsf{P}_{r}({\boldsymbol{x}}){\boldsymbol{h}}:=\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{2}}{\boldsymbol{h}}\quad\text{and}\quad\mathsf{P}_{\varphi}({\boldsymbol{x}}){\boldsymbol{h}}:=\left({\mathbf{I}}-\frac{{\boldsymbol{x}}{\boldsymbol{x}}^{\top}}{|{\boldsymbol{x}}|^{2}}\right){\boldsymbol{h}}.$ (2.14)
Abbreviating $\widehat{{\boldsymbol{s}}}:={{\boldsymbol{x}}}/{|{\boldsymbol{x}}|}$ (i.e., the radial unit vector) and $\rho:=|{\boldsymbol{x}}|$, we therefore have
$\mathsf{D}\boldsymbol{\Psi}({\boldsymbol{x}}){\boldsymbol{h}}=\widehat{{\boldsymbol{s}}}\,\partial_{1}\psi(\rho,\widehat{{\boldsymbol{s}}})\widehat{{\boldsymbol{s}}}^{\top}\mathsf{P}_{r}({\boldsymbol{x}}){\boldsymbol{h}}+\frac{\widehat{{\boldsymbol{s}}}}{\rho}\,\partial_{2}\psi(\rho,\widehat{{\boldsymbol{s}}})\mathsf{P}_{\varphi}({\boldsymbol{x}}){\boldsymbol{h}}+\frac{1}{\rho}\,\psi(\rho,\widehat{{\boldsymbol{s}}})\mathsf{P}_{\varphi}({\boldsymbol{x}}){\boldsymbol{h}}.$
The result then follows since the projections $\mathsf{P}_{r}({\boldsymbol{x}})$ and $\mathsf{P}_{\varphi}({\boldsymbol{x}})$ split the perturbation vector ${\boldsymbol{h}}$ into its radial and angular components and $\psi(\rho,\widehat{{\boldsymbol{s}}})=\rho+\chi(\rho)r(\widehat{{\boldsymbol{s}}})$.∎
Lemma 2.3 immediately implies that for $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in B_{2}$
$\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ $=\left(1+\chi^{\prime}(\widehat{\rho})r(\widehat{{\boldsymbol{s}}})\right)\left(1+\tfrac{\chi(\widehat{\rho})}{\widehat{\rho}}r(\widehat{{\boldsymbol{s}}})\right)^{d-1},$ (2.15a)
$\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})^{-1}$ $=\frac{1}{\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})}\begin{bmatrix}1+\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}r(\widehat{{\boldsymbol{s}}})&-\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}\dfrac{dr}{d\widehat{{\boldsymbol{s}}}}(\widehat{{\boldsymbol{s}}})\\ \mathbf{0}&\left(1+\chi^{\prime}(\widehat{\rho})r(\widehat{{\boldsymbol{s}}})\right)\,{\mathbf{I}}_{d-1}\end{bmatrix}\;.$ (2.15b)
By (2.4c), the bounds for $r(\widehat{{\boldsymbol{s}}})$, $-1/3\leq r(\widehat{{\boldsymbol{s}}})\leq 1/3$ for $\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1}$ (by Assumption 2.1), imply
$\dfrac{5}{3}\geq|1+\chi^{\prime}(\widehat{\rho})r(\widehat{{\boldsymbol{s}}})|\geq\dfrac{1}{3}\quad\text{and}\quad\dfrac{5}{3}\geq\Big{|}1+\frac{\chi(\widehat{\rho})}{\widehat{\rho}}r(\widehat{{\boldsymbol{s}}})\Big{|}\geq\dfrac{1}{3}\quad\text{ for all }\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1},0\leq\widehat{\rho}\leq 2;$
which, in turn, imply
$(5/3)^{d}\geq|\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r,\widehat{{\boldsymbol{x}}})|\geq 3^{-d}\quad\text{ for all}r\in{\cal R}\text{ and }\widehat{{\boldsymbol{x}}}\in B_{2},$ (2.16)
and, by estimating the Euclidean matrix norm via the Frobenius norm, for all $\widehat{{\boldsymbol{x}}}\in B_{2}$, $r\in\mathcal{R}$
$\left\|{\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r,\widehat{{\boldsymbol{x}}})}\right\|_{2}^{2}\leq\tfrac{5}{3}d+2\left\|{\frac{dr}{ds}(\widehat{{\boldsymbol{s}}})}\right\|_{2}^{2}\;,\quad\left\|{\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r,\widehat{{\boldsymbol{x}}})^{-1}}\right\|_{2}^{2}\leq{3^{d}}\left(\tfrac{5}{3}d+2\left\|{\frac{dr}{ds}(\widehat{{\boldsymbol{s}}})}\right\|_{2}^{2}\right).$ (2.17)
The following result – equivalence of norms of functions under the transformation $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ – is a consequence of (2.16) and (2.17). For a proof, see [41, Lemma 3.4] and [58, Appendix E].
Lemma 2.4 (Transformation of norms).
For $r\in{\cal R}$ define the pullback $\boldsymbol{\Phi}(r)^{*}v$ as $(\boldsymbol{\Phi}(r)^{*}v)(\widehat{{\boldsymbol{x}}}):=v(\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}}))$ for a function $v:B_{2}\to\mathbb{R}$.
  1. (i)
    If ${v\in L^{2}({B_{2}})}$, then ${\boldsymbol{\Phi}(r)^{*}v\in L^{2}({B_{2}})}$ and
    $\left(\tfrac{3}{5}\right)^{\frac{d}{2}}\left\|{v}\right\|_{L^{2}({B_{2}})}\leq\left\|{\boldsymbol{\Phi}(r)^{*}v}\right\|_{L^{2}({B_{2}})}\leq 3^{d/2}\left\|{v}\right\|_{L^{2}({B_{2}})}\;.$ (2.18)
  2. (ii)
    If $v\in H^{1}(B_{2})$, then $\boldsymbol{\Phi}(r)^{*}v\in H^{1}(B_{2})$ and
    $\left(\tfrac{3}{5}\right)^{\frac{d}{2}}\frac{1}{3^{d}\left(\tfrac{5}{3}d+2\left\|{\tfrac{dr}{ds}}\right\|_{C^{0}(\mathbb{S}^{d-1})}^{2}\right)}{\left\|{v}\right\|}_{H^{1}(B_{2})}$ $\leq{\left\|{\boldsymbol{\Phi}(r)^{*}v}\right\|}_{H^{1}(B_{2})}$
    $\leq 3^{d/2}\left(\tfrac{5}{3}d+2\left\|{\tfrac{dr}{ds}}\right\|_{C^{0}(\mathbb{S}^{d-1})}^{2}\right){\left\|{v}\right\|}_{H^{1}(B_{2})}\;.$ (2.19)
  3. (iii)
    If $r\in C^{2}(\mathbb{S}^{d-1})\cap{\cal R}$ and $v\in H^{2}(B_{2})$, then $\boldsymbol{\Phi}(r)^{*}v\in H^{2}(B_{2})$ and there exists $C_{2}>0$ depending on $d$ and $\left\|{r}\right\|_{C^{2}(\mathbb{S}^{d-1})}$ such that
    $C_{2}^{-1}{\left\|{v}\right\|}_{H^{2}(B_{2})}\leq{\left\|{\boldsymbol{\Phi}(r)^{*}v}\right\|}_{H^{2}(B_{2})}\leq C_{2}{\left\|{v}\right\|}_{H^{2}(B_{2})}\;.$ (2.20)
3 $k$-Explicit Norm Bounds for the Solution of the Transmission Problem
Recall from Section 1.4 that we are particularly interested in understanding how the domain of analyticity of $r\mapsto\widehat{u}(r)$, with $\widehat{u}(r)$ the solution of (2.8), behaves as the wavenumber $k$ becomes large. The key tools are the results of [53] concerning the stability of the transmission problem in variational formulation (2.7). These results are stated in the following $k$-scaled norms
$\left\|{u}\right\|^{2}_{H^{1}_{k}(B_{2})}:=k^{-2}\left\|{\nabla u}\right\|^{2}_{L^{2}(B_{2})}+\left\|{u}\right\|^{2}_{L^{2}(B_{2})}\;,\quad\left\|{L}\right\|_{(H^{1}_{k}(B_{2}))^{*}}:=\sup_{v\in H^{1}(B_{2})\setminus\{0\}}\frac{|L(v)|}{\left\|{v}\right\|_{H^{1}_{k}(B_{2})}}\;,$ (3.1)
and we record for later the definition that
$\left\|{u}\right\|^{2}_{H^{2}_{k}(B_{2})}:=k^{-4}|u|_{H^{2}(B_{2})}^{2}+k^{-2}\left\|{\nabla u}\right\|^{2}_{L^{2}(B_{2})}+\left\|{u}\right\|^{2}_{L^{2}(B_{2})}\;.$ (3.2)
As discussed in §1.3-1.4, the behaviour of the norm of the solution operator of the transmission problem with respect to $k$ depends on whether $n_{i}<1$ or $n_{i}>1$; if $n_{i}<1$ the norm of the solution operator from $L^{2}(B_{2})\to L^{2}(B_{2})$ grows like $k$ (see (3.3) below); if $n_{i}>1$ the solution operator can grow exponentially through $k_{j}$ for a discrete sequence $0<k_{1}<k_{2}<\ldots\infty$ [57], [62], [53, §6], although the growth for “most” $k$ is at most algebraic in $k$ [46]. We now recall bounds from [53, Theorem 3.1] on the solution when $n_{i}<1$.
Theorem 3.1 ($k$-explicit bound on the solution of the transmission problem from [53]).
Let $D$ be a star-shaped Lipschitz domain, and let $10<n_{i}<$ 1.
(i) If $L(v):=\int_{B_{2}}f\,\overline{v}\,{\rm d}S$ with $f\in L^{2}(B_{2})$, then the solution $u$ of the variational problem (2.7) satisfies
$\left\|{u}\right\|_{H^{1}_{k}(B_{2})}\leq{C_{{\rm sol,}1}(k,n_{i})}\left\|{f}\right\|_{L^{2}(B_{2})}\;,$ (3.3)
where
$C_{{\rm sol,}1}(k,n_{i}):=\frac{4k}{\sqrt{n_{i}}}\sqrt{1+\frac{1}{n_{i}}\left(1+\frac{d-1}{4k}\right)^{2}}.$ (3.4)
(ii) Given $L\in(H^{1}(B_{2}))^{*}$, the solution $u$ of the variational problem (2.7) satisfies
$\left\|{u}\right\|_{H^{1}_{k}(B_{2})}\leq C_{{\rm sol,}2}(k,n_{i})\left\|{L}\right\|_{(H^{1}_{k}(B_{2}))^{*}},$ (3.5)
where
$C_{{\rm sol,}2}(k,n_{i}):=\frac{1}{n_{i}}\Big{(}1+2C_{{\rm sol,}1}(k,n_{i})\Big{)}.$ (3.6)
The key points of this result for the present study are that (i) as $k\to\infty$, $C_{{\rm sol,}1}$ and $C_{{\rm sol,}2}$ both grow like $k$, and (ii) the ($k$-explicit) constants $C_{{\rm sol,}1}$ and $C_{{\rm sol,}2}$ are independent of $D$; i.e., the bounds (3.3) and (3.5) hold uniformly across all star-shaped Lipschitz domains.
References for the proof.
The bound (3.3) is proved in [53, Theorem 3.1]; indeed, (3.3) follows from [53, Equation 3.2] by (i) choosing $a_{i}=a_{o}=A_{D}=A_{N}=n_{o}=1$ and (ii) by bounding $k^{-2}\|\nabla u\|^{2}_{L^{2}(B_{2})}+n_{i}\|u\|_{L^{2}(B_{2})}^{2}$ below by $n_{i}(k^{-2}\|\nabla u\|^{2}_{L^{2}(B_{2})}+\|u\|_{L^{2}(B_{2})}^{2})$, and (iii) noting that, in the proof of [53, Theorem 3.1], ${\rm diam}(\Omega_{i})$ can be replaced by $2$ (when going from [53, Equation 5.3] to [53, Equation 5.4]). Having proved the bound (3.3), the bound (3.5) follows by, e.g., [36, Theorem 5.1]. ∎
Remark 3.2 (The Helmholtz exterior Dirichlet problem).
The analogous $k$-explicit bounds on the solution of the Helmholtz exterior Dirichlet problem that hold uniformly for all Lipschitz star-shaped $D$ were proved in [11, Corollary 3.9] (see also [36, Theorem 2.5]). The appropriate analogues of the results in the next section about $k$-explicit holomorphic dependence of the solution on the radial displacement function could therefore also be obtained for the Helmholtz exterior Dirichlet problem using these bounds in place of Theorem 3.1.
4 $k$-Explicit Holomorphic Dependence of the Solution on the Radial Displacement Function
For an approximation error analysis of polynomial surrogate models we invoke techniques from complex analysis. This requires extending the mapping $r\in{\cal R}\mapsto\widehat{u}(r)$, defined as the solution of the variational problem (2.8), to complex-valued radial displacement functions $r$, for which we adopt the notation $\mathfrak{r}$. The definition of the set ${\cal R}$ from Assumption 2.1 straightforwardly extends to $\mathbb{C}$-valued $r$. Then we write
:=$\mathfrak{R}:=\Big{\{}\mathfrak{r}\in C^{1}(\mathbb{S}^{d-1},\mathbb{C}):\,\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})}\leq\frac{1}{3}\Big{\}}\;.$ (4.1)
As explained in Section 1.2 our key task is to establish the holomorphy of $\mathfrak{r}\mapsto\hat{u}(\mathfrak{r})\in H^{1}(B_{R})$ and identify its domain of analyticity and its dependence on $k$. From now on we assume that $k\geq 1/2$ to avoid degeneracy of the constants in the bounds of Theorem 3.1.
For $\mathfrak{r}\in\mathfrak{R}$ let $\widehat{u}(\mathfrak{r})$ be the solution of the following variational problem: given $\widehat{L}\in(H^{1}(B_{2}))^{*}$,
$\mbox{find}\;\widehat{u}(\mathfrak{r})\in H^{1}(B_{2})\;\mbox{such that}\quad\widehat{a}(\mathfrak{r};\widehat{u},\widehat{v})=\widehat{L}(\widehat{v})\quad\mbox{for all}\;\;\widehat{v}\in H^{1}(B_{2})\;.$ (4.2)
Here, for $\mathfrak{r}\in\mathfrak{R}$, $\widehat{a}(\mathfrak{r};\cdot,\cdot)$ denotes the sesquilinear form from (2.8), (2.9), i.e.
$\widehat{a}(\mathfrak{r};\widehat{u},\widehat{v}):=&\int\limits_{B_{2}}{k^{-2}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}}(\widehat{{\boldsymbol{x}}})-\widehat{n}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\overline{\widehat{v}}(\widehat{{\boldsymbol{x}}})}\,{\mathrm{d}\widehat{{\boldsymbol{x}}}}-\int\limits_{\partial B_{2}}k^{-1}\mathsf{DtN}(\left.\mathstrut\widehat{u}(\mathfrak{r};\cdot)\right|_{\partial B_{2}})\overline{\widehat{v}}\,\mathrm{d}S,\end{aligned}$ (4.3)
with (now complex-valued) coefficients $\widehat{{\mathbf{A}}}(\mathfrak{r},\cdot)$ and $\widehat{n}(\mathfrak{r},\cdot)$, still given by the same formulae (2.9). The main result of this section is the following.
Theorem 4.1 ($k$-explicit condition for existence of $\widehat{u}(\mathfrak{r};\cdot)$ and related bounds).
(i) In the setting of §2, fix $n_{i}<1$, $d$, and, in addition, a constant $C_{\operatorname{Re}}\geq 2$.
Then there exist $C_{\operatorname{Im}},C_{1}>0,k_{0}\geq\frac{1}{2}$ such that for all $k\geq k_{0}$, for any $\widehat{L}\in(H^{1}_{k}(B_{2}))^{*}$, and for any
$\mathfrak{r}\in\mathfrak{R}\quad\text{with}\quad\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}\quad\text{and}\quad k\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Im}},$ (4.4)
the solution $\widehat{u}(\mathfrak{r};\cdot)$ of (4.2) exists, is unique, and satisfies the bound
$\left\|{\widehat{u}(\mathfrak{r};\cdot)}\right\|_{H^{1}_{k}(B_{2})}\leq C_{1}k\big{\|}\widehat{L}\big{\|}_{(H^{1}_{k}(B_{2}))^{*}}\;.$ (4.5)
(ii) If, in addition,
$\mathfrak{r}\in C^{2}(\mathbb{S}^{d-1})\quad\text{and}\quad\|\mathfrak{r}\|_{C^{2}(\mathbb{S}^{d-1})}\leq C_{S}\quad\text{for some}\quad C_{S}>0,$
then there exists $C_{2}=C_{2}(C_{S})>0$ such that, if $\widehat{L}(\widehat{v}):=\int_{B_{2}}f\,\overline{\widehat{v}}$ for $f\in L^{2}(B_{2})$ and $k\geq k_{0}$, then
$\left\|{\widehat{u}(\mathfrak{r};\cdot)}\right\|_{H^{2}_{k}(B_{2})}\leq C_{2}k\big{\|}f\big{\|}_{L^{2}(B_{2})}.$ (4.6)
The $\mathfrak{r}$-dependent coefficients $\widehat{{\mathbf{A}}}(\mathfrak{r};\cdot)$ and $\widehat{n}(\mathfrak{r};\cdot)$ in the sesquilinear form on the left-hand side of (2.8)/(4.2) are holomorphic on $\mathfrak{R}$. Let $F(\mathfrak{r},\widehat{u}(\mathfrak{r}))$ be defined as the left-hand side of (4.2) minus the right-hand side, so that $F(\mathfrak{r},\widehat{u}(\mathfrak{r}))=0$. The analytic implicit function theorem, e.g. [23, Thm. 15.3], applied to $F$ and Part (i) of Theorem 4.1 then implies the following result.
Corollary 4.2 (Domain of analyticity of $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})$).
Fix $d$, $n_{i}<1$, and a constant $C_{\operatorname{Re}}\geq 2$. Then there exist $C_{\operatorname{Im}},k_{0}\geq\frac{1}{2}$ such that for all $k\geq k_{0}$, given $\widehat{L}\in(H^{1}_{k}(B_{2}))^{*}$, the mapping
[…]
is analytic on the set
${\cal A}={\cal A}(k;C_{\operatorname{Re}},C_{\operatorname{Im}}):=\left\{\mathfrak{r}\in\mathfrak{R}:\;\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq{C_{\operatorname{Re}}},\;k\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Im}}\right\}.$ (4.7)
Remark 4.3.
(Analyticity of $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})\in H^{2}$) By Part (ii) of Theorem 4.1, an analogous result holds for the mapping $\mathfrak{r}\in C^{2}(\mathbb{S}^{d-1},\mathbb{C})\mapsto\widehat{u}(\mathfrak{r})\in H^{2}(B_{2})$ when $\widehat{L}(\widehat{v}):=\int_{B_{2}}f\,\overline{\widehat{v}}$ for $f\in L^{2}(B_{2})$. Although we do not use this result in the rest of the paper, additional smoothness of the target space of the analytic map $\mathfrak{r}\mapsto\widehat{u}(\mathfrak{r})$, translates into higher Sobolev regularity of the (analytic continuation of) the parametric solution. The higher Sobolev regularity of the analytic continuation of the parametric solution (as opposed to mere boundedness of this continuation in the energy norm of the PDE) is key in the convergence analysis of multilevel QMC methods; see, e.g., [24, §4.3.1] and the further discussion in §9.
To prove Theorem 4.1, we first record some results about the sesquilinear form in (4.2) for (complex valued) $\mathfrak{r}\in\mathfrak{R}$, as opposed to (real valued) $r\in\mathcal{R}$.
Lemma 4.4.
Fix $d\in\{2,3\}$ and $C_{\operatorname{Re}}\geq 2$. Then there exist $c_{1},c_{2}>0$ such that, if
$\mathfrak{r}\in\mathfrak{R}\quad\text{with}\quad\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}\quad\text{and}\quad\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq c_{1},$
then
$\operatorname{Re}\big{(}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\xi,\xi\big{)}_{2}\geq c_{2}\|\xi\|^{2}_{2}\quad\text{ for all }\xi\in\mathbb{R}^{d},\;\widehat{{\boldsymbol{x}}}\in B_{2}.$ (4.8)
By, e.g., [50, Page 122], the condition (4.8) ensures that the differential operator underlying the variational formulation (4.2) is strongly elliptic. We note that (4.8) is analogous to the condition [41, Equation 5.16] in [41, Assumption 5.11].
Proof of Lemma 4.4.
By the definition of $\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ (2.9a),
$\big{(}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\xi,\xi\big{)}_{2}=(\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}}))\|\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{-\top}\xi\|_{2}^{2};$
thus it is sufficient to bound $\|\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\top}\|_{2}$ from above and $\operatorname{Re}\big{(}\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{)}$ from below. The bound on $\|\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\top}\|_{2}$ from above follows by replacing $r$ by $\mathfrak{r}$ in the expression (2.10) for $\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ and then recognising that the bound (2.17) on $\|\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$ holds also for $\|\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\top}\|_{2}$. Finally, by the expression (2.15a) (with $r$ replaced by $\mathfrak{r}$) and the bound (2.16), $\operatorname{Re}(\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}}))$ is continuous with respect to $\operatorname{Im}\mathfrak{r}$ and bounded below by $3^{-d}$ when $\operatorname{Im}\mathfrak{r}=0$. Therefore there exist $c_{1}^{\prime},c_{2}^{\prime}>0$ such that
$\operatorname{Re}(\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}}))\geq c_{2}^{\prime}\quad\text{ when }\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq c_{1}^{\prime};$
the result then follows. ∎
We also need upper bounds for the transformed coefficient functions, which are supplied by the next lemma.
Lemma 4.5.
  1. (i)
    There exists ${C>0}$ depending only on ${d\in\{2,3\}}$ such that
    $\sup_{\widehat{{\boldsymbol{x}}}\in B_{2}}\big{\|}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{\|}_{2}\leq C\left(1+\|\mathfrak{r}\|_{C^{1}(\mathbb{S}^{d-1})}\right)^{2}\quad\text{and}\quad\sup_{\widehat{{\boldsymbol{x}}}\in B_{2}}\big{\|}\widehat{n}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{\|}_{2}\leq\left(1+2\left\|\mathfrak{r}\right\|_{C^{0}(\mathbb{S}^{d-1})}\right)^{d}.$ (4.9)
  2. (ii)
    Fix $d\in\{2,3\}$ and $c_{3}^{\prime}>0$ and assume that
    $\mathfrak{r}\in\mathfrak{R}\quad\text{ with }\quad\|\mathfrak{r}\|_{C^{2}(\mathbb{S}^{d-1})}\leq c_{3}^{\prime}\;.$ (4.10)
    Then there exists $c_{4}^{\prime}>0$ such that
    $\sup_{\widehat{{\boldsymbol{x}}}}\big{\|}\widehat{\boldsymbol{\nabla}}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{\|}_{2}\leq c_{4}^{\prime}\;\mbox{ for all }\;\widehat{{\boldsymbol{x}}}\in B_{2}\;.$
Lemma 4.4 and Part (i) of Lemma 4.5 have the following immediate corollary.
Corollary 4.6 (Continuity and Gårding inequality for the sesquilinear form).
Let $\widehat{a}(\mathfrak{r};\cdot,\cdot)$ denote the sesquilinear form on the left-hand side of (4.2). We fix $d\in\{2,3\}$ and $C_{\operatorname{Re}}>2$. Then there exists $c_{1}>0$ such that if
$\mathfrak{r}\in\mathfrak{R}\quad\text{with}\quad\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}\quad\text{and}\quad\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq c_{1},$
then there exist ($k$-independent) constants $c_{2},c_{5},c_{6}>0$ with
$|\widehat{a}(\mathfrak{r};\widehat{u},\widehat{v})|\leq c_{5}\big{\|}\widehat{u}\big{\|}_{H^{1}_{k}(B_{2})}\big{\|}\widehat{v}\big{\|}_{H^{1}_{k}(B_{2})}\quad\forall\widehat{u},\widehat{v}\in H^{1}(B_{2})$
and
$\operatorname{Re}\big{(}\widehat{a}(\mathfrak{r};\widehat{v},\widehat{v})\big{)}\geq c_{2}\big{\|}\widehat{v}\big{\|}^{2}_{H^{1}_{k}(B_{2})}-c_{6}\big{\|}\widehat{v}\big{\|}^{2}_{L^{2}(B_{2})}\quad\forall\widehat{v}\in H^{1}(B_{2}),$
and for all $k\geq\frac{1}{2}$.
Proof of Lemma 4.5.
By the explicit formula (2.15b),
$\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{-1}=\frac{\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}}{\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})}\;,\quad\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}:=\begin{bmatrix}1+\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}\mathfrak{r}(\widehat{{\boldsymbol{s}}})&-\frac{\chi(\widehat{\rho})}{\widehat{\rho}}\dfrac{d\mathfrak{r}}{d{\boldsymbol{s}}}(\widehat{{\boldsymbol{s}}})\\ \mathbf{0}&\left(1+\chi^{\prime}(\widehat{\rho})\mathfrak{r}(\widehat{{\boldsymbol{s}}})\right)\,{\mathbf{I}}_{d-1}\end{bmatrix},$ (4.11)
so that $\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ defined by (2.9a) can be written as
$\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})=\frac{1}{\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})}\,\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}\big{(}\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}\big{)}^{\top},\;\;\widehat{{\boldsymbol{x}}}\in B_{2}\;.$ (4.12)
From the expression (4.11) for $\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}$, the bounds on $\chi$ in (2.4) imply that there exists $C>0$ (depending only on $d$) such that
$\big{\|}\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}\big{\|}_{2}\leq C\big{(}1+\left\|{\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\big{)}\quad\text{ for all }\mathfrak{r}\in\mathfrak{R},\;\widehat{{\boldsymbol{x}}}\in B_{2}\;;$ (4.13)
a similar bound holds for $\|(\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp})^{\top}\|_{2}$.
We next observe that the lower bound (2.16) on the modulus of the determinant holds even for complex-valued $\mathfrak{r}\in\mathfrak{R}$, i.e.,
$|\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})|\geq 3^{-d}\;\text{ for all }\mathfrak{r}\in\mathfrak{R}\;,\;\;\widehat{{\boldsymbol{x}}}\in B_{2}\;.$ (4.14)
This is because $\left|1+\beta\mathfrak{r}(\widehat{{\boldsymbol{s}}})\right|\geq\left|1+\beta\operatorname{Re}\mathfrak{r}(\widehat{{\boldsymbol{s}}})\right|\geq\frac{1}{3}$ if $0\leq\beta\leq 2$ and $-\frac{1}{3}\leq\operatorname{Re}\mathfrak{r}(\widehat{{\boldsymbol{s}}})\leq\frac{1}{3}$, regardless of the imaginary part of $\mathfrak{r}(\widehat{{\boldsymbol{s}}})$. This completes the proof of the bound on $\sup_{\widehat{{\boldsymbol{x}}}}\big{\|}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{\|}_{2}$.
To prove the bound for $\widehat{n}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ we simply appeal to its definition (2.9b) and the expression (2.15a) for $\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ The proof of the bound on $\sup_{\widehat{{\boldsymbol{x}}}}\big{\|}\widehat{\boldsymbol{\nabla}}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{\|}_{2}$ in Part (ii) is similar, but more involved because of the differentiation (and thus requiring one more derivative of $\mathfrak{r}$ – contrast (4.9) and (4.10)). We omit the details. ∎
We are now in a position to prove Theorem 4.1. A central issue here is that for complex-valued $\mathfrak{r}\in\mathfrak{R}$ Theorem 3.1 is not immediately applicable. The crucial idea of the proof is to rewrite the variational problem (4.2) as
$\int_{B_{2}}\Big{(}k^{-2}\widehat{{\mathbf{A}}}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}}(\widehat{{\boldsymbol{x}}})-\widehat{n}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\overline{\widehat{v}}(\widehat{{\boldsymbol{x}}})\Big{)}{\rm d}\widehat{{\boldsymbol{x}}}-\int\nolimits_{\partial B_{2}}k^{-1}\mathsf{DtN}(\left.\mathstrut\widehat{u}(\mathfrak{r};\cdot)\right|_{\partial B_{2}})\overline{\widehat{v}}\,\mathrm{d}S$
$=\widehat{L}(\widehat{v})+\int_{B_{2}}\bigg{[}k^{-2}\Big{(}\widehat{{\mathbf{A}}}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\Big{)}\widehat{\boldsymbol{\nabla}}\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}(\widehat{{\boldsymbol{x}}})}-\Big{(}\widehat{n}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\widehat{n}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\Big{)}\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\overline{\widehat{v}}(\widehat{{\boldsymbol{x}}})\bigg{]}{\rm d}\widehat{{\boldsymbol{x}}}\;.$
We then transform back to the physical domain via $\widehat{{\boldsymbol{x}}}\mapsto{\boldsymbol{x}}:=\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})$ and obtain that $u(\mathfrak{r};\cdot):=\widehat{u}(\mathfrak{r},\boldsymbol{\Phi}^{-1}(\operatorname{Re}\mathfrak{r},{\boldsymbol{x}}))$ solves the variational problem:
find $u(\mathfrak{r};\cdot)\in H^{1}(B_{2})$ such that, for all $v\in H^{1}(B_{2})$,
$\int_{B_{2}}\Big{(}k^{-2}\boldsymbol{\nabla}u(\mathfrak{r};{\boldsymbol{x}})\cdot\boldsymbol{\nabla}v({\boldsymbol{x}})-n(\operatorname{Re}\mathfrak{r};{\boldsymbol{x}})u(\mathfrak{r};{\boldsymbol{x}})\,\overline{v}({\boldsymbol{x}})\Big{)}{\rm d}{\boldsymbol{x}}-\int_{\partial B_{2}}k^{-1}\mathsf{DtN}(u(\mathfrak{r};\cdot)|_{\partial B_{2}})\overline{v}\,{\rm d}S\\ =L(v)+\widetilde{L}(\mathfrak{r};v),$ (4.15)
where
$\widetilde{L}(\mathfrak{r};v)$ $:=\int_{B_{2}}\bigg{[}k^{-2}\widetilde{{\mathbf{A}}}(\mathfrak{r};{\boldsymbol{x}})\boldsymbol{\nabla}u(\mathfrak{r};{\boldsymbol{x}})\cdot\boldsymbol{\nabla}v({\boldsymbol{x}})-\widetilde{n}(\mathfrak{r};{\boldsymbol{x}})u(\mathfrak{r};{\boldsymbol{x}})\,\overline{v}({\boldsymbol{x}})\bigg{]}{\rm d}{\boldsymbol{x}}\;,$ (4.16)
$\widetilde{{\mathbf{A}}}(\mathfrak{r};{\boldsymbol{x}})$ $:=\left.\frac{\Big{(}\widehat{{\mathbf{A}}}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\Big{)}}{\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})}\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})\,\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\top}\right|_{\widehat{{\boldsymbol{x}}}\mapsto{\boldsymbol{x}}:=\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})}\;,$ (4.17)
$\widetilde{n}(\mathfrak{r};{\boldsymbol{x}})$ $:=\left.\frac{\Big{(}\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\Big{)}n_{0}(\widehat{{\boldsymbol{x}}})}{\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})}\right|_{\widehat{{\boldsymbol{x}}}\mapsto{\boldsymbol{x}}:=\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})}\;.$ (4.18)
The key point is that (4.15) is a variational formulation corresponding to a Helmholtz problem with real coefficients, and thus we can apply Theorem 3.1.
Lemma 4.7.
Given $d>0$, there exist $C_{\rm diff,1},C_{\rm diff,2}>0$ such that, for all $\mathfrak{r}\in\mathfrak{R}$,
$\left\|{\widetilde{n}(\mathfrak{r};\cdot)}\right\|_{L^{\infty}(B_{2})}$ $\leq C_{\rm diff,1}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})}\big{(}1+\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})}\big{)}^{d-1},$ (4.19a)
$\big{\|}\widetilde{{\mathbf{A}}}(\mathfrak{r},\cdot)\big{\|}_{L^{\infty}(B_{2})}$ $\leq C_{\rm diff,2}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\big{(}1+\left\|{\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\big{)}^{{2d-2}}.$ (4.19b)
Proof of Theorem 4.1 assuming Lemma 4.7.
(i) Given $C_{\rm Re}>0$, let $c_{1},c_{2}>0$ be as in Lemma 4.4. Assume that $\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}$ and $\|\operatorname{Im}\mathfrak{r}\|_{C^{1}(\mathbb{S}^{d-1})}\leq c_{1}$ (in the course of the proof we will restrict $\|\operatorname{Im}\mathfrak{r}\|_{C^{1}(\mathbb{S}^{d-1})}$ further). In particular, $\left\|{\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}+c_{1}=:C_{0}$.
By Corollary 4.6, the sesquilinear form $\widehat{a}(\cdot,\cdot)$ is continuous and satisfies a Gårding inequality, with constants independent of $\mathfrak{r}$. Therefore, by Fredholm theory, if, under the assumption of existence, one has a bound on the solution of the Helmholtz transmission problem in terms of the data $\widehat{L}$, then the solution exists and is unique (for more details of this method of arguing applied to the Helmholtz equation, see, e.g., [36, Lemma 3.5]).
It is therefore sufficient to assume that $\widehat{u}(\mathfrak{r};\cdot)$ exists and show that the bound (4.5) holds under the conditions on $\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}$ and $\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}$ that are implied by $\mathfrak{r}\in{\cal A}(k;C_{\operatorname{Re}},C_{\operatorname{Im}})$, with the constant $c_{1}$ in Lemma 4.4 given by $c_{1}=C_{\operatorname{Im}}/k$.
By the definition of $\widetilde{L}$ (4.16), the bounds (4.19), the assumption $\left\|{\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{0}$, and the Cauchy-Schwarz inequality
$|\widetilde{L}(v)|$ $\leq k^{-2}\big{\|}\widetilde{{\mathbf{A}}}(\mathfrak{r},\cdot)\big{\|}_{L^{\infty}(B_{2})}\left\|{\boldsymbol{\nabla}u(\mathfrak{r};\cdot)}\right\|_{L^{2}(B_{2})}\left\|{\boldsymbol{\nabla}v}\right\|_{L^{2}(B_{2})}+\left\|{\widetilde{n}(\mathfrak{r};\cdot)}\right\|_{L^{\infty}(B_{2})}\left\|{u(\mathfrak{r};\cdot)}\right\|_{L^{2}(B_{2})}\left\|{v}\right\|_{L^{2}(B_{2})}$
$\leq C_{\rm diff}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\Big{(}k^{-2}\left\|{\boldsymbol{\nabla}u(\mathfrak{r};\cdot)}\right\|_{L^{2}(B_{2})}\left\|{\boldsymbol{\nabla}v}\right\|_{L^{2}(B_{2})}+\left\|{u(\mathfrak{r};\cdot)}\right\|_{L^{2}(B_{2})}\left\|{v}\right\|_{L^{2}(B_{2})}\Big{)},$
with $C_{\rm diff}>0$ depending only on $C_{0},$ $C_{2}$, and on the constants in Lemma 4.7. Thus, by the definitions of $\|\cdot\|_{(H^{1}_{k}(B_{2}))^{*}}$ and $\|\cdot\|_{H^{1}_{k}(B_{2})}$ in (3.1) and the Cauchy–Schwarz inequality,
$\big{\|}\widetilde{L}\big{\|}_{(H^{1}_{k}(B_{2}))^{*}}$ $\leq C_{\rm diff}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\left\|{u(\mathfrak{r};\cdot)}\right\|_{H^{1}_{k}(B_{2})}.$
Applying the a priori bound (3.5), we obtain that the solution $u$ of (4.15) satisfies
$\left\|{u(\mathfrak{r};\cdot)}\right\|_{H^{1}_{k}(B_{2})}\leq C_{{\rm sol,}2}(k,n_{i})\Big{(}C_{\rm diff}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\left\|{u(\mathfrak{r};\cdot)}\right\|_{H^{1}_{k}(B_{2})}+\left\|{L}\right\|_{(H^{1}_{k}(B_{2}))^{*}}\Big{)}.$
Now, if
$C_{{\rm sol,}2}(k,n_{i})C_{\rm diff}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq 1/2,$ (4.20)
then
$\left\|{u(\mathfrak{r};\cdot)}\right\|_{H^{1}_{k}(B_{2})}\leq 2C_{{\rm sol,}2}(k,n_{i})\left\|{L}\right\|_{(H^{1}_{k}(B_{2}))^{*}}\;.$ (4.21)
If $k\geq 1/2$, then the explicit expressions (3.4) and (3.6) for $C_{{\rm sol,}1}$ and $C_{{\rm sol,}2}$, respectively, imply that
$C_{{\rm sol,}2}(k,n_{i})\leq\frac{k}{n_{i}}\left(1+\frac{{8}}{\sqrt{n_{i}}}\sqrt{1+\frac{1}{n_{i}}\left(1+\frac{d-1}{2}\right)^{2}}\right).$
We now convert the bound (4.21) on $u$ in terms of $L$ into the bound (4.5) on $\widehat{u}$ in terms of $\widehat{L}$. Indeed, (4.5) follows (for a suitable $C_{1}$ depending on $d$, $n_{i}$, and $C_{0}$) by using the expression for $C_{{\rm sol,}2}(k,2,n_{i})$ along with the norm equivalence from Lemma 2.4 combined with the bound $\left\|{\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\leq C_{0}$ (and observing that the norm equivalence in $\|\cdot\|_{H^{1}_{k}(B_{2})}$ implies a similar equivalence in $\|\cdot\|_{(H^{1}_{k}(B_{2}))^{*}}$). We now let
$C_{\operatorname{Im}}:=\left(\frac{2}{n_{i}}\left(1+\frac{8}{\sqrt{n_{i}}}\sqrt{1+\frac{1}{n_{i}}\left(1+\frac{d-1}{2}\right)^{2}}\right)C_{\rm diff}\right)^{-1}$
and $k_{0}:=\min\{C_{\operatorname{Im}}/c_{1},1/2\}$, so that the condition on $\operatorname{Im}\mathfrak{r}$ in (4.4) implies that $\operatorname{Im}\mathfrak{r}$ satisfies both (4.20) and $\|\operatorname{Im}\mathfrak{r}\|_{C^{1}(\mathbb{S}^{d-1})}\leq c_{1}$.
(ii) The assumption that $\widehat{L}(\widehat{v})=\int_{B_{2}}f\overline{\widehat{v}}$ for $f\in L^{2}(B_{2})$ implies that the variational problem (4.2) is equivalent to the (variational formulation of the) boundary-value problem
$k^{-2}\widehat{\boldsymbol{\nabla}}\cdot\big{(}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{)}$ $=-\widehat{n}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})-f(\widehat{{\boldsymbol{x}}})\quad\text{ for }\widehat{{\boldsymbol{x}}}\in B_{2},$
$k^{-1}\partial_{n}\widehat{u}$ $=\mathsf{DtN}\big{(}\widehat{u}|_{\partial B_{2}}\big{)}\quad\text{ on }\partial B_{2}.$
Then, by the $H^{2}$ regularity result of [47, Theorem 6.1], there exists $C>0$ (depending on $\sup_{\widehat{{\boldsymbol{x}}}}\|\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$, $\sup_{\widehat{{\boldsymbol{x}}}}\|\widehat{\boldsymbol{\nabla}}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$, and the constant $c_{2}$ in (4.8)) such that, for all $k\geq 1/2$, say,
$\big{\|}\widehat{u}\big{\|}_{H^{2}_{k}(B_{2})}\leq C\big{(}\big{\|}\widehat{u}\big{\|}_{H^{1}_{k}(B_{2})}+\big{\|}f\big{\|}_{L^{2}(B_{2})}\big{)}.$
The bound (4.6) then follows from the bound (4.5) from Part (i) and Lemmas 4.4 and 4.5 (with the assumption $\|\mathfrak{r}\|_{C^{2}(\mathbb{S}^{d-1})}\leq C$ required to apply Part (ii) of Lemma 4.5). ∎
It therefore remains to prove Lemma 4.7. To proceed, we record the following two elementary results.
Lemma 4.8.
Let $V$ be a normed vector space, $D\subset V$ some subset. If the matrix-valued functions ${\mathbf{F}}_{j}:D\to\mathbb{R}^{m\times m}$, $j=1,\ldots,n$, $m,n\in\mathbb{N}$, are such that there exist $C_{1},C_{2}>0$ such that, for all $n\in\mathbb{N}$ and for all $j=1,\ldots,n$,
$\left\|{{\mathbf{F}}_{j}({\mathbf{y}})-{\mathbf{F}}_{j}({\mathbf{z}})}\right\|\leq C_{1}\left\|{{\mathbf{y}}-{\mathbf{z}}}\right\|_{V}\;\forall{\mathbf{y}},{\mathbf{z}}\in D\quad\text{ and }\quad\left\|{{\mathbf{F}}_{j}({\mathbf{y}})}\right\|\leq C_{2}\;\forall{\mathbf{y}}\in D$ (4.22)
then
$\left\|{\prod_{j=1}^{n}{\mathbf{F}}_{j}({\mathbf{y}})-\prod_{j=1}^{n}{\mathbf{F}}_{j}({\mathbf{z}})}\right\|\leq nC_{1}C_{2}^{n-1}\left\|{{\mathbf{y}}-{\mathbf{z}}}\right\|_{V},$ (4.23)
[…]
Proof.
Defining a product to be ${\mathbf{I}}$ in case the lower index exceeds the upper, we have
$\prod_{j=1}^{n}{\mathbf{F}}_{j}({\mathbf{y}})-\prod_{j=1}^{n}{\mathbf{F}}_{j}({\mathbf{z}})$ $=\sum\limits_{k=1}^{n}\left(\prod\limits_{j=1}^{n-k+1}{\mathbf{F}}_{j}({\mathbf{y}})\cdot\prod\limits_{j=n-k+2}^{n}{\mathbf{F}}_{j}({\mathbf{z}})-\prod\limits_{j=1}^{n-k}{\mathbf{F}}_{j}({\mathbf{y}})\cdot\prod\limits_{j=n-k+1}^{n}{\mathbf{F}}_{j}({\mathbf{z}})\right)$
$=\sum\limits_{k=1}^{n}\left(\left(\prod\limits_{j=1}^{n-k}{\mathbf{F}}_{j}({\mathbf{y}})\right)({\mathbf{F}}_{n-k+1}({\mathbf{y}})-{\mathbf{F}}_{n-k+1}({\mathbf{z}}))\left(\prod\limits_{j=n-k+2}^{n}{\mathbf{F}}_{j}({\mathbf{z}})\right)\right)$
(where the right-hand side of the first equality is a telescoping sum). The result then follows from sub-multiplicativity and the triangle inequality. ∎
Lemma 4.9.
Given $z_{1},z_{2}\in\mathbb{C}$ with $|z_{1}|,|z_{2}|\geq L_{0}{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{>}}0$, for any two elements ${\mathbf{x}}_{1},{\mathbf{x}}_{2}\in V$ of a normed vector space $V$,
$\left\|\frac{{\mathbf{x}}_{1}}{z_{1}}-\frac{{\mathbf{x}}_{2}}{z_{2}}\right\|_{V}\leq L_{0}^{-2}\Big{(}\left\|{{\mathbf{x}}_{1}}\right\|_{V}\left|z_{2}-z_{1}\right|+|z_{1}|\left\|{{\mathbf{x}}_{2}-{\mathbf{x}}_{1}}\right\|_{V}\Big{)}\;.$ (4.24)
Proof.
The estimate is an immediate consequence of the triangle inequality and
$\frac{{\mathbf{x}}_{1}}{z_{1}}-\frac{{\mathbf{x}}_{2}}{z_{2}}=\frac{z_{2}{\mathbf{x}}_{1}-z_{1}{\mathbf{x}}_{2}}{z_{1}z_{2}}=\frac{(z_{2}-z_{1}){\mathbf{x}}_{1}+z_{1}({\mathbf{x}}_{1}-{\mathbf{x}}_{2})}{z_{1}z_{2}}.$
Proof of Lemma 4.7.
We first prove the bound on $\widetilde{n}$ in (4.19). By the bound (4.14), it is sufficient to bound the numerator in (4.18). We do this for fixed $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in B_{2}$ using the formula (2.15a) for $\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ and Lemma 4.8 with $V=C^{0}(\mathbb{S}^{d-1},\mathbb{C})$, $n=d$, $m=1$, $f_{1}(\mathfrak{r};\widehat{{\boldsymbol{x}}})=1+\chi^{\prime}(\widehat{\rho})\mathfrak{r}(\widehat{{\boldsymbol{s}}})$, $f_{\ell}(\mathfrak{r};\widehat{{\boldsymbol{x}}})=1+\frac{\chi(\widehat{\rho})}{\widehat{\rho}}\mathfrak{r}(\widehat{{\boldsymbol{s}}})$, and $\ell=2,\ldots,d$. With (2.4), we find the crude bounds
$|f_{\ell}(\mathfrak{r};\widehat{{\boldsymbol{x}}})|&\leq{1+2\left\|{\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})}},\\ |f_{\ell}(\mathfrak{r};\widehat{{\boldsymbol{x}}})-f_{\ell}(\mathfrak{r}^{\prime};\widehat{{\boldsymbol{x}}})|&\leq 2\left\|{\mathfrak{r}-\mathfrak{r}^{\prime}}\right\|_{C^{0}(\mathbb{S}^{d-1})}\end{aligned}\quad\text{ for all }\mathfrak{r}\in\mathfrak{R},\;\ell=1,\ldots,d\;,$ (4.25)
and for all $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in B_{2}$, $\widehat{{\boldsymbol{s}}}\in\mathbb{S}^{d-1}$. Since the bounds are uniform in $\widehat{{\boldsymbol{x}}}$ we conclude that
$\sup_{\widehat{{\boldsymbol{x}}}}\big{|}\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{|}\leq 2d\big{(}1+2\left\|{\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})})\big{)}^{d-1}\,\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})}\;.$ (4.26)
Combining this bound with (4.14) and using the fact that the real part of $\mathfrak{r}$ is bounded by $\frac{1}{3}$, we obtain (4.19a) with $C_{\rm diff,1}>0$ depending only on $d$.
We now prove the bound (4.19b) on $\widetilde{{\mathbf{A}}}(\mathfrak{r},{\boldsymbol{x}})$ (4.17). As above we fix $\widehat{{\boldsymbol{x}}}=\widehat{\rho}\widehat{{\boldsymbol{s}}}\in B_{2}$. Thanks to the lower bound for $|\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})|$ (4.14), it remains to bound (i) $\left\|{\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})}\right\|_{2}$ and (ii) $\|\widehat{{\mathbf{A}}}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\|_{2}$.
Regarding (i), in an almost-identical way to how we obtained (4.13), the explicit formula (2.10) and the bounds from (2.4) imply that there exists $C>0$ (depending only on $d$) such that
$\left\|{\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})}\right\|_{2}\leq C\big{(}1+\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\big{)}\quad\text{ for all }\mathfrak{r}\in\mathfrak{R},$
Regarding (ii), to bound $\widehat{{\mathbf{A}}}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})-\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ we use the expression (4.12) for $\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ and Lemma 4.8 with $m=d$, $n=3$, $V=C^{1}(\mathbb{S}^{d-1},\mathbb{C})$ and the terms ${\mathbf{F}}_{j}$ given by the factors in (4.12); i.e.
${\mathbf{F}}_{1}(\mathfrak{r}):=\big{(}\det\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{)}^{-1}{\mathbf{I}}\;,\quad{\mathbf{F}}_{2}(\mathfrak{r}):=\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}\;,\quad\text{ and }\quad{\mathbf{F}}_{3}(\mathfrak{r}):=\big{(}\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}\big{)}^{\top}\;.$ (4.27)
For all $\mathfrak{r}\in\mathfrak{R}$ and $\widehat{{\boldsymbol{x}}}\in B_{2}$, $\|{\mathbf{F}}_{1}(\mathfrak{r})\|_{2}\leq 3^{d}$ by (4.14), and
$\|{\mathbf{F}}_{1}(\operatorname{Re}\mathfrak{r})-{\mathbf{F}}_{1}(\mathfrak{r})\|_{2}\leq 9^{d}2d\big{(}1+2\|\mathfrak{r}\|_{C^{0}(\mathbb{S}^{d-1})}\big{)}^{d-1}\|\operatorname{Im}\mathfrak{r}\|_{C^{0}(\mathbb{S}^{d-1})}.$
by (4.26). Furthermore, $\|{\mathbf{F}}_{2}(\mathfrak{r})\|_{2}\leq C(1+\|\mathfrak{r}\|_{C^{1}(\mathbb{S}^{d-1})})$ by (4.13). Now, by the second equation in (4.11),
$\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}-\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}=-{\rm i}\begin{bmatrix}\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}\operatorname{Im}\mathfrak{r}(\widehat{{\boldsymbol{s}}})&-\dfrac{\chi(\widehat{\rho})}{\widehat{\rho}}\operatorname{Im}\dfrac{d\mathfrak{r}}{d{\boldsymbol{s}}}(\widehat{{\boldsymbol{s}}})\\ \mathbf{0}&\chi^{\prime}(\widehat{{\boldsymbol{s}}})\operatorname{Im}\mathfrak{r}(\widehat{{\boldsymbol{s}}}){\mathbf{I}}_{d-1}\;,\end{bmatrix},$
which implies that
$\|{\mathbf{F}}_{2}(\operatorname{Re}\mathfrak{r})-{\mathbf{F}}_{2}(\mathfrak{r})\|_{2}=\left\|{\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\operatorname{Re}\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}-\mathsf{D}_{\widehat{{\boldsymbol{x}}}}\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})^{\sharp}}\right\|_{2}\leq C\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\;,$
with a constant $C$ depending on $d$ alone. Finally, ${\mathbf{F}}_{3}$ satisfies exactly the same bounds as ${\mathbf{F}}_{2}$; thus we can apply Lemma 4.8 with
$C_{1}=C(1+\left\|{\mathfrak{r}}\right\|_{C^{0}(\mathbb{S}^{d-1})})^{d-1}\left\|{\operatorname{Im}\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\quad\text{and }\quad C_{2}=1+\left\|{\mathfrak{r}}\right\|_{C^{1}(\mathbb{S}^{d-1})}\;,$
to obtain (4.19b). ∎
5 FE Discretization: $k$-Explicit Error Bounds
In §4 we proved existence, uniqueness, and $k$-explicit bounds on the solution $\widehat{u}(\mathfrak{r};\cdot)$ to the variational problem (4.2), both for general data in $(H_{k}^{1}(B_{2}))^{*}$ and for $L^{2}$ data, see Theorem 4.1.
In this section we focus on computing approximations to $\widehat{u}(\mathfrak{r};\cdot)$, considering both general data in $(H_{k}^{1}(B_{2}))^{*}$ and the special case of $L^{2}$ data; the relevance of the latter is that the plane-wave scattering problem (1.1) can be reformulated as the solution of a problem with $L^{2}$ data – see §6.2 below.
Since the Dirichlet-to-Neumann map $\mathsf{DtN}$ is expensive to compute, we follow the commonly-used practice of approximation the radiation condition by truncation of the unbounded exterior domain by a radial perfectly-matched layer (PML).
5.1 Definition of radial PML truncation.
The solution $\widehat{u}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ to the variational problem (4.2) is approximated by the solution of the following problem. Let $2<R_{1}<R_{\rm tr}<\infty$ and let $\Omega_{\rm tr}\supset B(0,R_{\rm tr})$ be a bounded Lipschitz domain, and let $\widehat{u}_{\rm PML}\in H^{1}_{0}(\Omega_{\rm tr})$ be the solution of the variational problem
$\int_{\Omega_{\rm tr}}\bigg{(}k^{-2}\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\cdot\overline{\widehat{\boldsymbol{\nabla}}\widehat{v}(\mathfrak{r};\widehat{{\boldsymbol{x}}})}-\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{u}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\overline{\widehat{v}(\mathfrak{r};\widehat{{\boldsymbol{x}}})}\bigg{)}{\rm d}\widehat{{\boldsymbol{x}}}=\widehat{L}(\widehat{v})$ (5.1)
for all $\widehat{v}\in H^{1}_{0}(\Omega_{\rm tr})$, where $\widehat{{\mathbf{A}}}_{\rm PML}$ and $\widehat{n}_{\rm PML}$ are defined in terms of $\widehat{{\mathbf{A}}},\widehat{n}$, and functions $\alpha,\beta$ by
$\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}}):=\begin{cases}\widehat{{\mathbf{A}}}(\mathfrak{r};\widehat{{\boldsymbol{x}}})&\text{ for }|\widehat{{\boldsymbol{x}}}|\leq R_{1}\\ {\mathbf{H}}{\mathbf{D}}{\mathbf{H}}^{T}&\text{ for }|\widehat{{\boldsymbol{x}}}|>R_{1},\end{cases}\quad\text{ and }\quad\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}}):=\begin{cases}\widehat{n}(\mathfrak{r};\widehat{{\boldsymbol{x}}})&\text{ for }|\widehat{{\boldsymbol{x}}}|\leq R_{1}\\ \alpha(\widehat{\rho})\beta(\widehat{\rho})^{d-1}&\text{ for }|\widehat{{\boldsymbol{x}}}|>R_{1},\end{cases}$ (5.2)
where, in polar/spherical coordinates (with, as in §2, $|\widehat{{\boldsymbol{x}}}|=\widehat{\rho}$),
[…]
and
[…]
The functions $\alpha(\widehat{\rho})$ and $\beta(\widehat{\rho})$ are defined as follows: given a radial function $\widetilde{\sigma}$ and $R_{2}>R_{1}$ such that
$\widetilde{\sigma}(\widehat{\rho})=0\text{ for }\widehat{\rho}\leq R_{1},$ (5.3a)
$\widetilde{\sigma}(\widehat{\rho})\text{ is increasing}\text{ for}R_{1}\leq\widehat{\rho}\leq R_{2},\text{ and }$ (5.3b)
$\widetilde{\sigma}(\widehat{\rho})=\sigma_{0}>0\text{ for }\widehat{\rho}\geq R_{2},$ (5.3c)
let
$\sigma(\widehat{\rho}):=\big{(}\widehat{\rho}\widetilde{\sigma}(\widehat{\rho})\big{)}^{\prime},\quad\alpha(\widehat{\rho}):=1+{\rm i}\sigma(\widehat{\rho}),\quad\text{ and }\quad\beta(\widehat{\rho}):=1+{\rm i}\widetilde{\sigma}(\widehat{\rho}).$ (5.4)
We note that ${\mathbf{D}}={\mathbf{I}}_{d-1}$ and $n=1$ when $\widehat{\rho}=R_{1}$ and thus $\widehat{{\mathbf{A}}}_{\rm PML}$ and $\widehat{n}_{\rm PML}$ are continuous at $r=R_{1}$. We also note that $R_{\rm tr}$ can be $<R_{2}$, i.e., we allow truncation before $\widetilde{\sigma}$ reaches $\sigma_{0}$.
Remark 5.1 (The definition of the PML scaling function).
Here we have followed, e.g., [4, §2], [31, §1.2] and, starting from $\widetilde{\sigma}$, defined $\sigma$ in terms of $\widetilde{\sigma}$. Alternatively, one can start from a non-decreasing function $\sigma$ and define $\widetilde{\sigma}$ such that the first equation in (5.4) holds; see, e.g., [20, §3], [48, §2], [42, §4], [12, §2]. The notation $\alpha$ and $\beta$ is also used by [48, 49].
5.2 Properties of the sesquilinear form and solution operator of the PML problem
Lemma 5.2 (Sign property of $\operatorname{Re}(\widehat{{\mathbf{A}}}_{\rm PML})$).
Given $C_{\operatorname{Re}}>0$, let ${\mathcal{A}},C_{\operatorname{Im}},k_{0}$ be as in Corollary 4.2. Given $\widetilde{\sigma}$ as in (5.3), there exists $C>0$ such that, for all $\mathfrak{r}\in{\mathcal{A}}$ as defined in (4.7), for all $\widehat{{\boldsymbol{x}}}\in\mathbb{R}^{d}$, and for all $\xi\in\mathbb{R}^{d}$,
$\operatorname{Re}\big{(}\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\xi,\xi\big{)}_{2}\geq C\|\xi\|^{2}_{2}.$ (5.5)
Proof.
This follows by combining Lemma 4.4 and the analogous result about the sign-definiteness of the PML coefficient, see, e.g., [32, Lemma 2.3] (where $\widetilde{\sigma}(r)=f_{\theta}(r)/r$ in the notation of [32, Lemma 2.3]). In using Lemma 4.4, we use the fact that the constants $C_{\operatorname{Im}},k_{0}$ from Corollary 4.2 are such that if $\mathfrak{r}\in\mathcal{A}$ defined by (4.7), then the assumptions of Lemma 4.4 are satisfied. ∎
Lemma 5.2 and Corollary 4.6 then imply the following analogue of Corollary 4.6.
Corollary 5.3.
The sesquilinear form in the variational problem (5.1) is continuous and satisfies a Gårding inequality, with both the continuity constant and the constants in the Gårding inequality independent of $\mathfrak{r}$ for $\mathfrak{r}\in{\mathcal{A}}$.
Theorem 5.4 (PML solution operator inherits behaviour of non-truncated solution operator).
Given $n_{i}<1,$ $d=2$ or $3$, and $C_{\operatorname{Re}}>0$, let ${\mathcal{A}}$ and $C_{\operatorname{Im}}$ be as in (4.7), Corollary 4.2. Then, given $\widetilde{\sigma}\in C^{3}(0,\infty)$ as in (5.3), there exists $C,k_{1}\geq\frac{1}{2}$ such that, for all $\mathfrak{r}\in{\mathcal{A}}$, for any $\widehat{L}\in(H^{1}_{k}(B_{2}))^{*}$, and for all $k\geq k_{1}$, the solution $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ of (5.1) exists, is unique, and satisfies the bound
$\left\|{\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)}\right\|_{H^{1}_{k}(\Omega_{\rm tr})}\leq Ck\big{\|}\widehat{L}\big{\|}_{(H^{1}_{k}(\Omega_{\rm tr}))^{*}}\;.$ (5.6)
Proof.
This follows from [31, Theorem 1.6], where $k_{1}$ is taken as the maximum of $k_{0}$ from Corollary 4.2 and the corresponding lower bound on $k$ coming from [31, Theorem 1.6].
Indeed, [31, Theorem 1.6] proves that the solution operator of the PML problem is bounded by the solution operator of the original (i.e., non-truncated) problem for all scattering problems in the black-box scattering framework of [60].
The variational problem (4.2) does not fit in the black-box framework as described in, e.g., [28, Chapter 4], since the PDE is not formally self-adjoint when $\operatorname{Im}\mathfrak{r}\neq 0$ (because the coefficients are complex). Nevertheless, the results from the black-box framework required to prove [31, Theorem 1.6] still hold. Indeed, [31, Theorem 1.6] follows from [31, Lemma 3.3], and the two ingredients of this lemma that depend on the contents of the black box are (i) that the PML problem is Fredholm of index zero, and (ii) agreement away from the scaling region of the solution operators of the complex-scaled and unscaled Helmholtz problems (see, e.g., [28, Theorem 4.37]).
Regarding (i): for the variational problem (5.1), this property follows from Corollary 5.3 and, e.g., [50, Theorem 2.34]. Regarding (ii): inspecting the proof of [28, Theorem 4.37], we see that the only place where this proof uses that the contents of the black-box are self adjoint is in finding a complex $k$ where the (unscaled) problem can be shown to have a unique solution and satisfies the natural bound on the solution operator (in terms of $|k|$-dependence); see [28, Lemma 4.3]. In our case, if $k={\rm i}\lambda$, then one can prove directly from the variational formulation (using Lemma 5.2 and the definition of $\widehat{n}$ (2.9b)) that the problem (4.2) with $\mathfrak{r}\in{\mathcal{A}}$ has a unique solution and satisfies the bound (4.5) with $k$ replaced by $\lambda$. ∎
Remark 5.5.
As the transformations $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ leave the PML-zone invariant, $\mathfrak{r}\mapsto\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ is still analytic on ${\cal A}$.
5.3 The accuracy of PML truncation for $k$ large.
Theorem 5.6 (Radial PMLs are exponentially accurate for $k$ large).
Given $n_{i}<1$, $d=2$ or $3$, and $C_{\operatorname{Re}}>0$, let ${\mathcal{A}}$ and $C_{\operatorname{Im}}$ be as in Corollary 4.2. Given $\widetilde{\sigma}\in C^{3}(0,\infty)$ as in (5.3), and $\epsilon>0$, there exist $C_{\rm PML,1},C_{\rm PML,2},k_{1}>0$ such that the following is true for all $R_{\rm tr}>(1+\epsilon)R_{1}$, $\mathfrak{r}\in{\mathcal{A}}$, and $k\geq k_{1}$.
Given $\widehat{f}\in L^{2}(B_{2})$, let $\widehat{u}$ be the solution of the variational problem (4.2) with $\widehat{L}(\widehat{v})=\int_{B_{2}}\widehat{f}\overline{\widehat{v}}{\rm d}\widehat{{\boldsymbol{x}}}$, and let $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ be the solution of the variational problem (5.1) with the same $\widehat{L}$. Then $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ exists, is unique, and satisfies
$\|\widehat{u}(\mathfrak{r};\cdot)-\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)\|_{H^{1}_{k}(B(0,R_{1}))}\leq C_{\rm PML,1}\exp\Big{(}-C_{\rm PML,2}k\big{(}R_{\rm tr}-(1+\epsilon)R_{1}\big{)}\Big{)}\big{\|}\widehat{f}\big{\|}_{L^{2}(B_{2})}.$ (5.7)
References for the proof.
This result is a special case of [31, Theorem 1.4]. Just like [31, Theorem 1.6] (which we used in the proof of Theorem 5.4) the proof of [31, Theorem 1.4] uses agreement away from the scaling region of the solution operators of the complex-scaled and unscaled Helmholtz problems [28, Theorem 4.37]; this result holds for our problem as explained in the proof of Theorem 5.4. ∎
Theorem 5.6 shows that the error in approximating $\widehat{u}(\mathfrak{r};\cdot)$ by $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ decreases exponentially in the wavenumber $k$ and the PML width $R_{\rm tr}-R_{1}$, uniformly for $\mathfrak{r}\in{\mathcal{A}}$.
5.4 The accuracy of the $h$-FEM approximation of the PML solution
We consider subspaces $(V_{h})_{h>0}$ of $H^{1}(\Omega_{\rm tr})$ satisfying the following assumption.
Assumption 5.7.
For some “polynomial degree” $p\in\mathbb{N}$, the subspaces $(V_{h})_{h>0}$ satisfy the following: there exists a constant $C>0$ such that, for all $h>0$ and $0\leq\ell\leq p$, given $v\in H^{1}_{0}(\Omega_{\rm tr})\cap H^{\ell+1}(\Omega_{\rm tr})$ there exists $\mathcal{I}_{h,p}v\in V_{h,p}$ such that
$|v-\mathcal{I}_{h,p}v|_{H^{j}(\Omega_{\rm tr})}\leq Ch^{\ell+1-j}\|v\|_{H^{\ell+1}(\Omega_{\rm tr})}.$ (5.8)
Assumption 5.7 holds when $(V_{h})_{h>0}$ consists of functions that are continuous in $\Omega_{\rm tr}$ and piecewise polynomials of total degree $p$ on a shape-regular family of simplicial triangulations of $\Omega_{\rm tr}$, indexed by the meshwidth $h$; see, e.g., [18, Theorem 17.1], [5, Proposition 3.3.17].
Let $\widehat{a}_{\rm PML}(\mathfrak{r};\cdot,\cdot)$ denote the sesquilinear form on the left-hand side of (5.1). The sequence of Galerkin solutions of the PML problem is defined by
$\boxed{\text{find}\quad\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}\in V_{h}\,\,\text{ such that }\,\,\widehat{a}_{\rm PML}(\mathfrak{r};\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h},\widehat{v}_{h})=\widehat{L}(\widehat{v}_{h})\quad\text{ for all }\widehat{v}_{h}\in V_{h}\;.}$ (5.9)
Our result about $k$-explicit convergence of the $h$-FEM (Theorem 5.8 below) requires that, for $\gamma\geq p-1$, $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)\in C^{\gamma,1}$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\in C^{\gamma-1,1}$ with the corresponding norms bounded independently of $\mathfrak{r}$ (the reason for this is explained in the discussion preceding Lemma 5.10 and the proof of Theorem 5.8). For simplicity, we assume that $\gamma=p-1$, i.e., given the regularity of $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)$ and $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$, we take the minimal polynomial degree to obtain the best-possible FEM convergence result for that regularity.
By the definitions of $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)$ in terms of $\widehat{{\mathbf{A}}}(\mathfrak{r};\cdot)$ (2.9a), for $\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\cdot)\in C^{p-1,1}$ we need $\mathfrak{r}$-uniform control of the $C^{p,1}$ norm of $r(\mathfrak{r};\cdot)$. Similarly, by the definition of $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})$ in terms of $\widehat{n}(r;\widehat{{\boldsymbol{x}}})$ (2.9b), for $\widehat{n}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\in C^{p-2,1}$ we need $\mathfrak{r}$-uniform control of the $C^{p-1,1}$ norm of $r(\mathfrak{r};\cdot)$. We therefore consider the following subsets of ${\mathcal{A}}$ (4.7): for $p=1,2,3,...$, define  [^6] For simplicity we make the constant on the right-hand side of the bound on the $C^{p,1}$ norms in (5.10) $C_{\operatorname{Re}}$, but in principle this could be a different constant.
${\mathcal{A}}_{p}={\mathcal{A}}_{p}(k;C_{\operatorname{Re}},C_{\operatorname{Im}}):=\big{\{}\mathfrak{r}\in{\mathcal{A}}(k;C_{\operatorname{Re}},C_{\operatorname{Im}})\,:\,\left\|{\operatorname{Re}\mathfrak{r}}\right\|_{C^{p,1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}\big{\}},\;p\in\mathbb{N}.$ (5.10)
Theorem 5.8 ($k$-explicit quasioptimality of $h$-FEM, uniform for $\mathfrak{r}\in{\mathcal{A}}_{p}$).
Given $n_{i}<1$, $d=2$ or $3$, $C_{\operatorname{Re}}>0$, $p\in\mathbb{N}$, suppose that
  • both the PML scaling function $\widetilde{\sigma}$ and $\partial\Omega_{\rm tr}$ are $C^{p,1}$ regular,
  • $k_{1}$ is as in Theorem 5.4,
  • $(V_{h})_{h>0}$ is as in Assumption 5.7.
Then there exists $C_{\rm FEM,j},j=1,2,3,4$ such that, for all $\mathfrak{r}\in{\mathcal{A}}_{p}$ (as defined in (5.10)), for all $\widehat{L}\in(H^{1}_{k}(\Omega_{\rm tr}))^{*}$, and for all $k\geq k_{1}$, if $h$ is such that
${(hk)^{2p}}k\leq C_{\rm FEM,1},$ (5.11)
then the solution $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}$ of (5.9) exists, is unique, and satisfies
$\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}\leq C_{\rm FEM,2}{\Big{(}1+(hk)^{p}k\Big{)}}\min_{\widehat{v}_{h}\in V_{h}}\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{v}_{h}\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}$ (5.12)
and
$\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{L^{2}(\Omega_{\rm tr})}\leq C_{\rm FEM,3}{\Big{(}hk+(hk)^{p}k\Big{)}}\min_{\widehat{v}_{h}\in V_{h}}\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{v}_{h}\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}.$ (5.13)
Suppose, in addition, that
  • $\widehat{L}(\widehat{v})=\int_{\Omega_{\rm tr}}\widehat{f}\overline{\widehat{v}}{\rm d}\widehat{{\boldsymbol{x}}}$ with $\widehat{f}\in{H^{p}}(\Omega_{\rm tr})$ that satisfies for some $C_{f}>0$
    $|\widehat{f}|_{H^{\ell}(\Omega_{\rm tr})}\leq C_{f}k^{\ell}\|\widehat{f}\|_{L^{2}(\Omega_{\rm tr})}\,\,\text{ for all }\,\,\ell=0,\ldots,p-1\,\,\text{ and }\,\,k\geq k_{1}.$ (5.14)
Then there is a constant $C_{\rm FEM,4}>0$ (depending on $C_{f}$) such that, for all $\mathfrak{r}\in{\mathcal{A}}_{p}$,
$\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}\leq C_{\rm FEM,4}\Big{(}1+(hk)^{p}k\Big{)}(hk)^{p}\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}$ (5.15)
and
$\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{L^{2}(\Omega_{\rm tr})}\leq C_{\rm FEM,4}\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}.$ (5.16)
Remark 5.9 (Interpreting Theorem 5.8 and the mesh threshold (5.11)).
Theorem 5.8 implies that
  1. 1.
    If $(hk)^{p}k$ is sufficiently small, then the Galerkin solutions are quasioptimal, with constant in independent of $k$, as $k\to\infty$ (by (5.12)).
  2. 2.
    If $(hk)^{2p}k$ is sufficiently small and the data is “$k$-oscillatory” in the sense of (5.14), then the relative $H^{1}_{k}$ error is controllably small, uniformly in $k$, as $k\to\infty$ (by (5.15)).
These thresholds are observed empirically to be sharp (going back to the work of Ihlenburg and Babuška in 1-d [43, 44]); see the discussion in [33, §1.3] and the references therein.
In the limit $h\to 0$ for fixed $k$, the bounds (5.15) and (5.16) recover the well-known results that, for $f\in H^{p-1}$, the Galerkin error in $H^{1}$ is $O(h^{p})$ (5.15), and the Galerkin error in $L^{2}$ is $O(h^{p+1})$ (5.16).
A key ingredient in the proofs of Theorem 5.8 is the standard elliptic regularity shift property for the highest-order terms in the PDE. We record this result for our problem (i.e., the PML problem (5.1)), emphasizing that the constants in the bound are independent of $\mathfrak{r}$ for $\mathfrak{r}\in{\mathcal{A}}_{p}$.
Lemma 5.10 (Elliptic regularity shift property, uniform for $\mathfrak{r}\in{\mathcal{A}}_{p}$).
Suppose that $\widetilde{\sigma}$ and $\partial\Omega_{\rm tr}$ are both $C^{p,1}$, $p\in\mathbb{N}$. Given $d=2$ or $3$ and $C_{\operatorname{Re}}>0$, let $C_{\operatorname{Im}}$ be as in Corollary 4.2. Then there exists $C>0$ such that for all $\mathfrak{r}\in{\mathcal{A}}_{p}$ and $j=1,\ldots,p+1$, given $\widehat{f}\in H^{j-2}(\Omega_{\rm tr})$, the solution of
^$-\widehat{\boldsymbol{\nabla}}\cdot\big{(}\widehat{{\mathbf{A}}}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\widehat{\boldsymbol{\nabla}}\widehat{w}(\mathfrak{r};\widehat{{\boldsymbol{x}}})\big{)}=\widehat{f}(\widehat{{\boldsymbol{x}}})\quad\text{ for }\widehat{{\boldsymbol{x}}}\in\Omega_{\rm tr}\quad\text{ and }\quad\widehat{w}(\widehat{{\boldsymbol{x}}})=0\quad\text{ for }\widehat{{\boldsymbol{x}}}\in\partial\Omega_{\rm tr}$
satisfies
$|\widehat{w}|_{H^{j}(\Omega_{\rm tr})}\leq C\big{\|}\widehat{f}\big{\|}_{H^{j-2}(\Omega_{\rm tr})}.$ (5.17)
Furthermore, the analogous bound hold when $\widehat{{\mathbf{A}}}_{\rm PML}$ is replaced by either $\overline{(\widehat{{\mathbf{A}}}_{\rm PML})^{T}}$ or $\operatorname{Re}(\widehat{{\mathbf{A}}}_{\rm PML}):=(\widehat{{\mathbf{A}}}_{\rm PML}+\overline{(\widehat{{\mathbf{A}}}_{\rm PML})^{T}})/2$.
Proof.
We first claim that it is sufficient to prove that
$|\widehat{w}|_{H^{j}(\Omega_{\rm tr})}\leq C^{\prime}\big{(}\left\|{\widehat{w}}\right\|_{H^{1}(\Omega_{\rm tr})}+\big{\|}\widehat{f}\big{\|}_{H^{j-2}(\Omega_{\rm tr})}\big{)}$ (5.18)
for some $C^{\prime}$ independent of $\mathfrak{r}$. Indeed, Lemma 5.2, Part (i) of Lemma 4.5 and the Lax–Milgram theorem imply that $\|\widehat{w}\|_{H^{1}(\Omega_{\rm tr})}\leq C^{\prime\prime}\|\widehat{f}\|_{L^{2}(\Omega_{\rm tr})}$ with $C^{\prime\prime}$ independent of $\mathfrak{r}$, and combining this with (5.18) yields (5.17).
By, e.g., [50, Theorem 4.18] the bound (5.18) holds if (i) (5.5) holds, (ii) $\partial\Omega_{\rm tr}$ is $C^{p,1}$, and (iii) $\widehat{{\mathbf{A}}}_{\rm PML}\in C^{p-1,1}(\Omega_{\rm tr})$ with norm bounded independently of $\mathfrak{r}$. The property (i) holds by Lemma 5.2 and the property (ii) holds by assumption. For the property (iii), the fact that $\widetilde{\sigma}\in C^{p,1}$ implies that $\widehat{{\mathbf{A}}}_{\rm PML}$ (defined by (5.2)) restricted to the PML region $|\widehat{{\boldsymbol{x}}}|\geq R_{1}$ is $C^{p-1,1}$. Arguing as in Lemma 4.5, we have that if $\mathfrak{r}\in\mathcal{A}_{p}$ then $\widehat{{\mathbf{A}}}_{\rm PML}$ restricted to $|\widehat{{\boldsymbol{x}}}|\leq R_{1}$ is in $C^{p,1}$ with norm bounded independently of $\mathfrak{r}$. ∎
Proof of Theorem 5.8.
We obtain the bounds (5.12), (5.13), and (5.15) under the mesh threshold (5.11) by showing that the PML problem (5.1) and its Galerkin discretisation (5.9) fit into the class of problems considered by [33]. Analytic dependence of the Galerkin solution on $\mathfrak{r}$ then follows by noting that, by (5.12) and the triangle inequality, $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)_{h}$ satisfies the same bound in terms of the data as $\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)$ (i.e., the bound (5.6)) and then arguing as in the proof of Corollary 4.2 using the analytic implicit function theorem.
The paper [33] proves the bounds (5.12)-(5.16) for Helmholtz problems whose sesquilinear forms are continuous, satisfy a Gårding inequality, and satisfy elliptic regularity. We now show how Corollary 5.3 and Lemma 5.10 imply that the variational problem (5.1) fits into this framework, with constants independent of $\mathfrak{r}$ for $\mathfrak{r}\in{\mathcal{A}}_{p}$.
Corollary 5.3 immediately shows that [33, Equations 1.6 and 1.7] are satisfied (i.e., the sesquilinear form is continuous and satisfies a Gårding inequality). For the elliptic regularity assumptions of [33, Assumptions 1.2 and 1.6], we need that there exists $C>0$ such that, for $j=1,\ldots,p+1$ and $\mathfrak{r}\in{\mathcal{A}}_{p}$,
$\left\|{\widehat{w}}\right\|_{H^{j}_{k}(\Omega_{\rm tr})}\leq C\Big{(}\left\|{\widehat{w}}\right\|_{L^{2}}+\big{\|}k^{-2}\nabla\cdot(\widehat{{\mathbf{A}}}_{\rm PML}\nabla\widehat{w})+\widehat{n}_{\rm PML}\widehat{w}\big{\|}_{H^{j-2}_{k}(\Omega_{\rm tr})}\Big{)},$ (5.19)
as well as the analogous bound with $\widehat{{\mathbf{A}}}_{\rm PML}$ replaced by either $\overline{(\widehat{{\mathbf{A}}}_{\rm PML})^{T}}$ or $\operatorname{Re}(\widehat{{\mathbf{A}}}_{\rm PML}):=(\widehat{{\mathbf{A}}}_{\rm PML}+\overline{(\widehat{{\mathbf{A}}}_{\rm PML})^{T}})/2$. In (5.19) $H^{j}_{k}$ is the usual $H^{j}$ norm, but with each derivative weighted by $k^{-1}$, as in (3.1), (3.2). By multiplying (5.17) by $k^{-j}$, we obtain that there exists $C>0$ such that, for $j=1,\ldots,p+1$ and $\mathfrak{r}\in{\mathcal{A}}_{p}$,
$\left\|{\widehat{w}}\right\|_{H^{j}_{k}(\Omega_{\rm tr})}\leq C\big{\|}k^{-2}\nabla\cdot(\widehat{{\mathbf{A}}}_{\rm PML}\nabla\widehat{w})\big{\|}_{H^{j-2}_{k}(\Omega_{\rm tr})}.$
Now, since $\widehat{n}_{\rm PML}\in C^{p-1}$ with norm bounded independently of $\mathfrak{r}\in{\mathcal{A}}_{p}$, by, e.g., [37, Theorem 1.4.1.1, page 21] there exists $C>0$ such that
$\|\widehat{n}_{\rm PML}\widehat{w}\|_{H^{j-2}_{k}(\Omega_{\rm tr})}\leq C\|\widehat{w}\|_{H^{j-2}_{k}(\Omega_{\rm tr})}\quad\text{ for }j=1,\ldots,p+1\,\text{ and }\,\mathfrak{r}\in{\mathcal{A}}_{p}.$ (5.20)
Therefore, for $j=1,\ldots,p+1$ and $\mathfrak{r}\in{\mathcal{A}}_{p}$,
$\left\|{\widehat{w}}\right\|_{H^{j}_{k}(\Omega_{\rm tr})}\leq C\Big{(}\left\|{\widehat{w}}\right\|_{H^{j-2}_{k}(\Omega_{\rm tr})}+\big{\|}k^{-2}\nabla\cdot(\widehat{{\mathbf{A}}}_{\rm PML}\nabla\widehat{w})+\widehat{n}_{\rm PML}\widehat{w}\big{\|}_{H^{j-2}_{k}(\Omega_{\rm tr})}\Big{)},$
and the required bound (5.19) follows by induction.
Assumption 5.7 is [33, Assumption 4.8], and then (5.12)-(5.15) follow from [33, Equations 4.16, 4.17, and 4.19]. The bound (5.16) is not stated explicitly in [33], but follows from the displayed equation before [33, Remark 2.3] (one repeats the arguments that obtain (5.15) from (5.12), but now one starts from (5.13)). ∎
6 Quantity of Interest (QoI): Far-Field Pattern
6.1 Definition of and expressions for the far-field pattern
If $v$ is a solution of the Helmholtz equation $(-k^{-2}\Delta-1)v=0$ outside $B_{R_{0}}$ for some $R_{0}>0$ and $v$ satisfies the Sommerfeld radiation condition (2.6) (i.e., $v$ is outgoing), then the far-field pattern of $v$, $v_{\infty}:\mathbb{S}^{d-1}\to\mathbb{C}$, is defined by
$v_{\infty}({\boldsymbol{q}})=\lim_{\rho\to\infty}\Big{(}\rho^{(d-1)/2}\exp(-{\rm i}k\rho)\,v(\rho{\boldsymbol{q}})\Big{)}\;,\quad{\boldsymbol{q}}\in\mathbb{S}^{d-1}$ (6.1)
(this limit exists and is a smooth function of ${\boldsymbol{q}}$ by, e.g., [21, Corollary 3.7]).
The quantity of interest for our scattering problem (1.1) is the far field pattern $u_{\infty}:\mathbb{S}^{d-1}\to\mathbb{C}$ of the scattered wave $u^{\mathrm{scat}}:=u-u^{\mathrm{inc}}$.
We now give an expression for the far-field pattern of an outgoing Helmholtz solution from [54, Theorem 2.2] (written in a slightly-more general way here).
Theorem 6.1 (Expression for far-field pattern as integral over subset of domain).
Suppose that, for some $R_{0}>0$, $v\in C^{2}(\mathbb{R}^{d}\setminus B_{R_{0}})$ satisfies the Helmholtz equation $(-k^{-2}\Delta-1)v=0$ in $\mathbb{R}^{d}\setminus\overline{B_{R_{0}}}$ and the Sommerfeld radiation condition (2.6). Let $\psi\in C^{2}(\mathbb{R}^{d};[0,1])$ be such that $\psi\equiv 0$ in a neighbourhood of $B_{R_{0}}$ and $\psi\equiv 1$ on $(B_{R_{1}})^{c}$ for some $R_{1}>R_{0}$. Then
$v_{\infty}({\boldsymbol{q}})=C(d,k)\int_{\operatorname{supp}\nabla\psi}v({\boldsymbol{y}})\Big{(}\Delta\psi({\boldsymbol{y}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi({\boldsymbol{y}})\Big{)}\exp\big{(}-{\rm i}k{\boldsymbol{y}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}{\boldsymbol{y}},$ (6.2)
where
$C(2,k)=\frac{1}{\sqrt{k}}\frac{{\rm e}^{{\rm i}\pi/4}}{2\sqrt{2\pi}}\quad\text{ and }\quad C(3,k)=\frac{1}{4\pi}.$ (6.3)
Since the proof is relatively short, we give it here.
Proof of Theorem 6.1.
Let
$G({\boldsymbol{x}},{\boldsymbol{y}}):=\frac{{\rm i}}{4}H_{0}^{(1)}(k|{\boldsymbol{x}}-{\boldsymbol{y}}|)\,\,\text{ for }d=2,\quad\text{ and }\quad G({\boldsymbol{x}},{\boldsymbol{y}}):=\frac{\exp({\rm i}k|{\boldsymbol{x}}-{\boldsymbol{y}}|)}{4\pi|{\boldsymbol{x}}-{\boldsymbol{y}}|}\,\,\text{ for}d=3,$
and recall that $(\Delta_{{\boldsymbol{y}}}+k^{2})G({\boldsymbol{x}},{\boldsymbol{y}})=-\delta({\boldsymbol{x}}-{\boldsymbol{y}})$. Green’s integral representation states that if $w\in C^{2}(\overline{B_{R}})$ then, for ${\boldsymbol{x}}\in B_{R}$,
$w({\boldsymbol{x}})=\int_{\partial B_{R}}\left(\frac{\partial w}{\partial\nu}({\boldsymbol{y}})G({\boldsymbol{x}},{\boldsymbol{y}})-w({\boldsymbol{y}})\frac{\partial G({\boldsymbol{x}},{\boldsymbol{y}})}{\partial\nu({\boldsymbol{y}})}\right){\rm d}S({\boldsymbol{y}})-\int_{B_{R}}G({\boldsymbol{x}},{\boldsymbol{y}})(\Delta+k^{2})w({\boldsymbol{y}}){\rm d}{\boldsymbol{y}},$
where $\nu({\boldsymbol{y}}):={\boldsymbol{y}}/|{\boldsymbol{y}}|$; see, e.g., [21, Theorem 3.1]. Let $w=v\psi$ and observe that $w\in C^{2}(\overline{B_{R}})$ since $\psi\equiv 0$ on a neighbourhood of $B_{R_{0}}$. With this choice of $w$, as $R\to\infty$, the integral over $\partial B_{R}$ tends to zero, since both $v$ and $G({\boldsymbol{x}},\cdot)$ (for ${\boldsymbol{x}}$ fixed) satisfy the Sommerfeld radiation condition (see, e.g., [21, Last equation in the proof of Theorem 3.3]). Now,
$(\Delta+k^{2})w=(\Delta+k^{2})(\psi v)=2\nabla\psi\cdot\nabla v+v\Delta\psi+\psi(\Delta+k^{2})v=2\nabla\psi\cdot\nabla v+v\Delta\psi,$
since $(\Delta+k^{2})v=0$ on $\operatorname{supp}\psi$; therefore, if $|{\boldsymbol{x}}|>R_{1}$, then
$v({\boldsymbol{x}})=-\int_{\operatorname{supp}\nabla\psi}G({\boldsymbol{x}},{\boldsymbol{y}})\Big{(}2\nabla\psi\cdot\nabla v+v\Delta\psi\Big{)}({\boldsymbol{y}}){\rm d}{\boldsymbol{y}}.$
Letting $|{\boldsymbol{x}}|\to\infty$ and using (6.1) and (when $d=2$) the large-argument asymptotics of $H_{0}^{(1)}(\cdot)$ (see, e.g., [56, Equation 10.17.5]), we find that
$v_{\infty}({\boldsymbol{q}})=-C(d,k)\int_{\operatorname{supp}\nabla\psi}\Big{(}2\nabla\psi\cdot\nabla v+v\Delta\psi\Big{)}({\boldsymbol{y}})\exp\big{(}-{\rm i}k{\boldsymbol{y}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}{\boldsymbol{y}}.$
The result (6.2) then follows by integrating by parts (using the divergence theorem) the term involving $\nabla\psi\cdot\nabla v$ (moving the derivative from $v$ onto $\psi$). ∎
6.2 Formulating the solution of the plane-wave scattering problem as the outgoing solution of a Helmholtz problem with $L^{2}$ data
We now recall how to formulate the solution of the plane-wave scattering problem of (1.1) as the outgoing solution of a Helmholtz problem with source $f\in L^{2}(B_{2})$, so that it can be approximated by PML truncation (with the error then given by Theorem 5.6).
One option is to solve for the scattering field $u^{\mathrm{scat}}:=u-u^{\mathrm{inc}}$, which satisfies the Sommerfeld radiation condition (2.6) (by (1.1b)) and
$(-k^{-2}\Delta-n)u^{\mathrm{scat}}=-(1-n)u^{\mathrm{inc}};$ (6.4)
since the right-hand side of this PDE is compactly supported in $B_{2}$, PML truncation can be used to approximate $u^{\mathrm{scat}}$ (with the error then controlled by Theorem 5.6).
A second option is described in the following lemma. Although this second option is more complicated than the first, the second option has the advantage that, when $u^{\mathrm{inc}}$ is an incident plane wave, the $L^{2}$-norm of the right-hand side of the PDE behaves like $O(k^{-1})$ for $k\to\infty$ (see (6.6) below), whereas the right-hand side of the PDE (6.4) is uniformly bounded with respect to $k$. Recall that slower growth in $k$ of the right-hand side implies slower growth of the solution (by (5.6)) and thus stronger bounds on the finite-element solution (by (5.15)).
Lemma 6.2 (Transmission solution formulated as an outgoing Helmholtz solution with $L^{2}$ data).
Given $u^{\mathrm{inc}}$, let $u$ be the solution of the Helmholtz transmission problem (1.1) with $n$ given by (1.2) and $D$ as described in §2.1. Given $\eta>0$, let $\varphi\in C^{2}_{\rm comp}(\mathbb{R}^{d};[0,1])$ be such that
$\varphi\equiv 1\,\,\text{ on }B_{2-\eta}\quad\text{ and }\quad\varphi\equiv 0\,\,\text{ on }\,\,\mathbb{R}^{d}\setminus B_{2-\eta/2}.$ (6.5)
Let $u^{\mathrm{alt}}$ be the outgoing solution to
$(-k^{-2}\Delta-n)u^{\mathrm{alt}}=-k^{-2}\big{(}2\nabla\varphi\cdot\nabla u^{\mathrm{inc}}+u^{\mathrm{inc}}\Delta\varphi\big{)}=:f^{\rm alt}.$ (6.6)
Then
$u^{\mathrm{alt}}=\varphi u^{\mathrm{inc}}+(u-u^{\mathrm{inc}})=u-(1-\varphi)u^{\mathrm{inc}},$ (6.7)
and thus $u^{\mathrm{alt}}\equiv u$ on $B_{2-\eta}$ and $u^{\mathrm{alt}}=u^{\mathrm{scat}}$ on $\mathbb{R}^{d}\setminus B_{2-\eta/2}$.
Proof.
Given $\varphi$ as above, the function $u-(1-\varphi)u^{\mathrm{inc}}$ satisfies both the Sommerfeld radiation condition (2.6) and
$(-k^{-2}\Delta-n)\big{(}u-(1-\varphi)u^{\mathrm{inc}}\big{)}=-(-k^{-2}\Delta-n)(1-\varphi)u^{\mathrm{inc}}$ $=-(-k^{-2}\Delta-1)(1-\varphi)u^{\mathrm{inc}}$
$=-k^{-2}\big{(}2\nabla\varphi\cdot\nabla u^{\mathrm{inc}}+u^{\mathrm{inc}}\Delta\varphi\big{)},$
where we have used that $n\equiv 1$ on $\operatorname{supp}(1-\varphi)$ in the second equality. By the uniqueness of outgoing solutions of the transmission problem, (6.7) holds. ∎
We now combine Theorem 6.1 and Lemma 6.2 to give an expression for the far-field pattern of the scattered wave in the transmission problem (1.1) in terms of $u^{\mathrm{alt}}$.
Corollary 6.3 (Far-field pattern in terms of $u^{\mathrm{alt}}$ in the nominal domain).
Let $\lambda>0$ be as in §2, (2.4), and choose $\eta>0$ such that $\eta<\lambda$. Let further $\psi\in C^{2}(\mathbb{R}^{d};[0,1])$ be such that $\psi\equiv 0$ in a neighbourhood of $B_{2-\lambda}$ and $\psi\equiv 1$ in a neighbourhood of $\mathbb{R}^{d}\setminus B_{2-\eta}$. Let $\varphi\in C^{2}(\mathbb{R}^{d};[0,1])$ be as in (6.5), and let $u^{\mathrm{alt}}$ be the outgoing solution of (6.6). Then
$\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};{\boldsymbol{q}})$ $=C(d,k)\int\limits_{\operatorname{supp}\nabla\psi}\Big{(}\widehat{u}^{\rm alt}(\mathfrak{r};\widehat{{\boldsymbol{y}}})-\widehat{u}^{\rm inc}(\widehat{{\boldsymbol{y}}})\Big{)}\Big{(}\Delta\psi(\widehat{{\boldsymbol{y}}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi(\widehat{{\boldsymbol{y}}})\Big{)}\cdot$
$\hskip 227.62204pt\exp\big{(}-{\rm i}k\widehat{{\boldsymbol{y}}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}\widehat{{\boldsymbol{y}}}\;,\quad{\boldsymbol{q}}\in\mathbb{S}^{d-1}\;.$ (6.8)
Furthermore, given $C_{\operatorname{Re}}>2$, let ${\mathcal{A}}={\mathcal{A}}(k;C_{\operatorname{Re}},C_{1})$ for $k>\frac{1}{2}$ be as in (4.7) and Corollary 4.2.
Then the map ${\cal A}\to L^{2}({\mathbb{S}^{d-1}}):\mathfrak{r}\mapsto\widehat{u}^{\rm scat}_{\infty}$ is holomorphic, and there exist $C^{\prime},C^{\prime\prime}>0$ such that, for all $k\geq\frac{1}{2}$,
$\sup\limits_{\mathfrak{r}\in{\cal A}}\left\|{\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};\cdot)}\right\|_{L^{\infty}(\mathbb{S}^{d-1})}\leq C^{\prime}C(d,k)k\left\|{\widehat{u}^{\rm inc}}\right\|_{H^{1}_{k}(B_{2})}\leq C^{\prime\prime}C(d,k)k.$ (6.9)
The constant $C(d,k)$ is as in (6.3); in particular, as $k\to\infty$, $C(d,k)=O(k^{(d-3)/2})$.
In Corollary 6.3 we choose $\psi$ such that $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ is the identity on $\operatorname{supp}\psi$ (by (2.4)); this choice is not necessary, i.e., an analogous expression to (6.8) holds if $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$ is not the identity on $\operatorname{supp}\psi$, but this analogous expression is slightly more complicated than (6.8).
Proof of Corollary 6.3.
By Theorem 6.1, $u^{\mathrm{scat}}_{\infty}$ is given by the right-hand side of (6.2) with $v$ replaced by $u-u^{\mathrm{inc}}$. The definition of $\psi$ implies that $\operatorname{supp}\nabla\psi\subset B_{2-\eta}$, and Lemma 6.2 implies that $u\equiv u^{\mathrm{alt}}$ on $B_{2-\eta}$. Therefore, $u^{\mathrm{scat}}_{\infty}$ is given by the right-hand side of (6.2) with $v$ replaced by $u^{\mathrm{alt}}-u^{\mathrm{inc}}$.
We now map the integral in (6.2) back to the nominal domain using $\widehat{{\boldsymbol{x}}}\mapsto\boldsymbol{\Phi}(r;\widehat{{\boldsymbol{x}}})$; by (2.4). This transformation is the identity on $\mathbb{R}^{d}\setminus B_{2-\lambda}$, and thus on $\operatorname{supp}\nabla\psi$, and the expression (6.8) follows.
By Corollary 4.2, the map $\mathfrak{r}\mapsto\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)$ is holomorphic for $\mathfrak{r}\in\mathcal{A}$. Since the mapping $\widehat{u}^{\rm alt}\mapsto\widehat{u}_{\infty}^{\rm scat}$ is linear and all other terms entering the integrand in (6.8) are independent of $\mathfrak{r}$, the map ${\cal A}\to L^{2}({\mathbb{S}^{d-1}})$, $\mathfrak{r}\mapsto\widehat{u}_{\infty}$, is holomorphic. Finally, by (4.5), the definition (3.1) of $\|\cdot\|_{(H_{k}^{1}(B_{2}))^{*}}$, and by (6.6), for $k\geq k_{0}$ it holds
$\sup_{\mathfrak{r}\in\mathcal{A}}\big{\|}\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)\big{\|}_{H^{1}_{k}(B_{2})}\leq C_{2}k\big{\|}\widehat{f}^{\rm alt}\big{\|}_{L^{2}(B_{2})}\leq C^{\prime}\big{\|}\widehat{u}^{\rm inc}\big{\|}_{H^{1}_{k}(B_{2})}.$ (6.10)
The bound (6.9) then follows by combining (6.10) and (6.8). ∎
6.3 Accuracy of the PML $h$-Galerkin FEM approximation of the far-field pattern
We now use the results of §5 to bound the error in the far-field pattern when PML truncation and the $h$-FEM are used to compute an approximation to $\widehat{u}^{\rm alt}$, and this approximation used in the expression (6.8).
As above, let $\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)$ be the outgoing solution of (6.6) transformed to the nominal domain. Let $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)$ be the solution of the PML variational formulation (5.1) with the corresponding right-hand side. Let $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}$ be the corresponding solution of the Galerkin equations (5.9), and let
$\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}:=C(d,k)\int\limits_{\operatorname{supp}\nabla\psi}\begin{aligned} \Big{(}\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}-\widehat{u}^{\rm inc}(\widehat{{\boldsymbol{y}}})\Big{)}\Big{(}\Delta\psi(\widehat{{\boldsymbol{y}}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi(\widehat{{\boldsymbol{y}}})\Big{)}\cdot\\ \exp\big{(}-{\rm i}k\widehat{{\boldsymbol{y}}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}\widehat{{\boldsymbol{y}}}\end{aligned}$ (6.11)
(i.e., (6.8) with $\widehat{u}^{\rm alt}(\mathfrak{r};\widehat{{\boldsymbol{y}}})$ replaced by $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}$) denote the corresponding approximate farfield.
Corollary 6.4 (Error in the computed far-field pattern).
Under the assumptions of Theorem 5.8 and with $k_{1}$ as in Theorem 5.4, there exists $C_{\rm FF,1}$ such that for all $k\geq k_{1}$, for all $\mathfrak{r}\in{\mathcal{A}}_{p}$, and for all $h,k$, and $p$ satisfying (5.11),
$\left\|{\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};\cdot)-\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}}\right\|_{L^{\infty}(\mathbb{S}^{d-1})}$
$\qquad\leq\,C_{\rm FF,1}C(d,k)\Big{[}\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}k+\exp\Big{(}-C_{\rm PML,2}k\big{(}R_{\rm tr}-(1+\epsilon)R_{1}\big{)}\Big{)}\Big{]}\left\|{\widehat{u}^{\rm inc}}\right\|_{H^{1}_{k}(B_{2})},$ (6.12)
where $C(d,k)$ is given for $d=2,3$ by (6.3).
Since $\|\widehat{u}^{\rm inc}\|_{H^{1}_{k}(B_{2})}\leq C^{\prime}$, the bound (6.12) shows that the error in the far-field pattern (6.12) is controlled uniformly in $k$ provided that $(hk)^{p}k$ is sufficiently small. Since $C(2,k)=O(k^{-1/2})$, when $d=2$ the error in the far-field pattern decreases with increasing $k$ subject to this mesh threshold.
Proof of Corollary 6.4.
By (6.8) and (6.11),
$\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};{\boldsymbol{q}})-\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$
$=C(d,k)\int_{\operatorname{supp}\nabla\psi}\Big{(}\widehat{u}^{\rm alt}(\mathfrak{r};\widehat{{\boldsymbol{y}}})-\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\widehat{{\boldsymbol{y}}})_{h}\Big{)}\Big{(}\Delta\psi(\widehat{{\boldsymbol{y}}})-2{\rm i}k\,{\boldsymbol{q}}\cdot\nabla\psi(\widehat{{\boldsymbol{y}}})\Big{)}\exp\big{(}-{\rm i}k\widehat{{\boldsymbol{y}}}\cdot{\boldsymbol{q}}\big{)}\,{\rm d}\widehat{{\boldsymbol{y}}}.$
By the Cauchy–Schwarz inequality and the triangle inequality, for some $C^{\prime}>0$,
$\left\|{\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};\cdot)-\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}}\right\|_{L^{\infty}(\mathbb{S}^{d-1})}\leq kC(d,k)C^{\prime}\big{\|}\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)-\widehat{u}_{\rm PML}^{\rm alt}(\mathfrak{r};\cdot)_{h}\big{\|}_{L^{2}(B_{2})}\\ \leq kC(d,k)C^{\prime}\Big{[}\big{\|}\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)-\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)\big{\|}_{L^{2}(B_{2})}+\big{\|}\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{L^{2}(B_{2})}\Big{]}\;.$ (6.13)
To bound the second term in parentheses on the right-hand side of (6.13), we use the FE error bound (5.16) and then the bound (5.6) applied to $\|\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)\|_{H^{1}(\Omega_{\rm tr})}$ to obtain that
$\big{\|}\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{L^{2}(B_{2})}$ $\leq\big{\|}\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}\big{\|}_{L^{2}(\Omega_{\rm tr})}$
$\leq C_{\rm FEM,4}\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}\big{\|}\widehat{u}_{\rm PML}(\mathfrak{r};\cdot)\big{\|}_{H^{1}_{k}(\Omega_{\rm tr})}$
$\leq C_{\rm FEM,4}\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}Ck\big{\|}\widehat{f}^{\rm alt}\big{\|}_{L^{2}(B_{2})},$ (6.14)
where we have used that $\widehat{f}^{\rm alt}$ defined by (6.6) satisfies the $k$-oscillatory property in Theorem 5.8 (to apply (5.16)) and is supported in $B_{2}$ (so that the final norm in (6.14) is over $B_{2}$). To bound the first term in the parentheses on the right-hand side of (6.13), we use the PML error bound (5.7). Combining this with (6.13) and (6.14), we obtain that, for some $C^{\prime\prime}>0$,
$\left\|{\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};\cdot)-\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}}\right\|_{L^{\infty}(\mathbb{S}^{d-1})}$
$\qquad\leq kC(d,k)C^{\prime\prime}\Big{[}\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}k+\exp\Big{(}-C_{\rm PML,2}k\big{(}R_{\rm tr}-(1+\epsilon)R_{1}\big{)}\Big{)}\Big{]}\|\widehat{f}^{\rm alt}\|_{L^{2}(B_{2})}.$
The result (6.12) then follows by recalling from (6.10) that $\|\widehat{f}^{\rm alt}\|_{L^{2}(B_{2})}\leq Ck^{-1}\|\widehat{u}^{\rm inc}\|_{H^{1}_{k}(B_{2})}$. ∎
Remark 6.5 (Comparison of Theorem 5.8 with the results of [29]).
The following two steps in the proof of Corollary 6.4 might appear over-simplistic:
  1. 1.
    using bounds involving the $L^{2}(B_{2})$ norm of the data $f^{\rm alt}$, while for plane-wave scattering $f^{\rm alt}\in H^{s}(B_{2})$ for all $s>0$ (see (6.6)), and
  2. 2.
    estimating the integral in the expression (6.8) for the far-field pattern using the Cauchy–Schwarz inequality, instead of using, say, a duality argument.
The paper [29] both
  1. 1.
    bounds the Galerkin error in terms of $\|f\|_{H^{s}}$ for arbitrary $s>0$, and
  2. 2.
    bounds $L(u-u_{h})$, where $L(u)=\int u\overline{z}$ for $z\in H^{s^{\prime}}$, using a duality argument.
Nevertheless, for the particular case of data $f$ coming from a plane-wave (i.e., $f^{\rm alt}$), and the functional $L(\cdot)$ being the far-field pattern expressed as (6.8), the results of [29] do not give a better result than Corollary 6.4.
We now briefly justify this statement; strictly speaking, the results in [29] cover the $hp$-FEM applied to the constant-coefficient Helmholtz equation, but in principle they can be extended to the $h$-FEM (for arbitrary $p$) for the variable-coefficient Helmholtz equation using the ideas of [13] and [33]. In this discussion we are interested in these results applied to the FE error $\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)-\widehat{u}^{\rm alt}_{\rm PML}(\mathfrak{r};\cdot)_{h}$ (as in Corollary 6.4), but to lighten notation we just talk about $u-u_{h}$. Furthermore, in this discussion we assume that the norm of the solution operator scales as $\sim k$ (as in Theorem 3.1).
A standard duality argument (see, e.g., [29, Proposition 2.1]) shows that
$\left\|{u-u_{h}}\right\|_{L^{2}}\leq C\left\|{u-u_{h}}\right\|_{H^{1}_{k}}\sup_{0\neq g\in L^{2}}\min_{v_{h}\in V_{h}}\frac{\left\|{\mathcal{S}^{*}g-v_{h}}\right\|_{H^{1}_{k}}}{\left\|{g}\right\|_{L^{2}}},$ (6.15)
and [29, Lemma 2.3] uses similar ideas to prove that
$\left|\int(u-u_{h})z\right|\leq C\left\|{u-u_{h}}\right\|_{H^{1}_{k}}\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}},$ (6.16)
where $\mathcal{S}^{*}:L^{2}\to H^{1}_{k}$ is the Helmholtz adjoint solution operator. The work [29] focusses on these bounds when the Galerkin solution is quasi-optimal, in which case the $H^{1}_{k}$ errors are bounded, uniformly in $k$, by the best approximation error. Recall from Theorem 5.8 and Remark 5.9 that a sufficient (and empirically necessary) condition for quasioptimality is that $(hk)^{p}k$ be sufficiently small.
If $z\in L^{2}$, then
$\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}}\leq C\Big{(}hk+(hk)^{p}k\Big{)}\left\|{z}\right\|_{L^{2}}$ (6.17)
by [51, 52], [13, Lemma 2.13], [33, Theorem 1.7]. Furthermore, if $z\in H^{p-1}$ is $k$-oscillatory, in that it satisfies the bound (5.14) (with $\widehat{f}$ replaced by $z$), then
$\min_{v_{h}\in V_{h}}\left\|{\mathcal{S}^{*}z-v_{h}}\right\|_{H^{1}_{k}}\leq C(hk)^{p}k\left\|{z}\right\|_{L^{2}}$ (6.18)
by [29, Part (ii) of Theorem 4.4]. The bounds (6.17) and (6.18) also hold with $\mathcal{S}^{*}$ replaced by $\mathcal{S}$ since $\mathcal{S}^{*}f=\overline{\mathcal{S}\overline{f}}$.
Therefore, if $u$ is the solution of the Helmholtz equation with $k$-oscillatory (in the sense of (5.14)) right-hand side $f\in H^{p-1}$ and $h$ is such that $(hk)^{p}k\leq C$ for sufficiently-small $C>0$, then the combination of (6.15), quasioptimality, (6.17), and (6.18) implies that
$\left\|{u-u_{h}}\right\|_{L^{2}}\leq C\Big{(}hk+(hk)^{p}k\Big{)}(hk)^{p}k\left\|{f}\right\|_{L^{2}}.$ (6.19)
Furthermore, if $z\in H^{p-1}$ is $k$-oscillatory, as in the case of the far-field pattern (6.8), then the combination of (6.16), quasioptimality, and (6.18) implies that
$\left|\int(u-u_{h})z\right|\leq C(hk)^{p}k\left\|{f}\right\|_{L^{2}}(hk)^{p}k\left\|{z}\right\|_{L^{2}}.$ (6.20)
The bound (6.19) is the same as (6.14), and the bound (6.20) is no better than (6.19) when $\|z\|_{L^{2}}\sim 1$ and when $(hk)^{p}k$ is fixed (i.e., when one chooses the least restrictive condition on $h$ allowed by the theory as $k\to\infty$). That is, in this setting of data coming from a plane-wave with the quantity of interest being the far-field pattern expressed as the linear functional (6.8), the results of [29] indeed do not give a better result than Corollary 6.4. (Note that the bounds on the error in the functional in [29, Part (ii) of Corollary 4.5 and Part (ii) of Corollary 4.6] assume that $\|z\|_{H^{s^{\prime}}}$ is independent of $k$, and thus these results have better $k$-dependence than stated above.)
7 $k$- and $h$-Explicit, Parametric Holomorphy of $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$
We now study the parametric holomorphy of the PML-Galerkin approximated far-field pattern $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$ defined in (6.11).
To this end, we assume that the shape variation is modelled as in Section 2.1, with displacement functions $r\in{\cal R}$ satisfying Assumption 2.1. Shape uncertainty will be introduced by placing a (probability) measure on a suitable set of radial displacement functions: $r=r(\omega)$, $\omega\in\Omega$, with $\Omega$ the set of elementary events in a probability space. For purposes of computation, we assume affine-parametric dependence of $r$ on a sequence ${\mathbf{y}}:=\left(y_{j}\right)_{j\in\mathbb{N}}$ of parameters. As is customary in computational UQ for PDEs (see, e.g., [59]), the measure will be constructed as product of probability measures on the co-ordinates $y_{j}$. This is to say that the dependence of $r$ on $\omega$ is expressed through a sequence $\boldsymbol{Y}=\left(Y_{j}\right)_{j\in\mathbb{N}}$ of independent identically distributed (i.i.d), $(-1,1)$-valued random variables $Y_{j}=Y_{j}(\omega)$.
We thus introduce an affine uncertainty-parametrization of the scatterer geometry with globally supported basis functions $r_{j}({\boldsymbol{s}})$, $j\in\mathbb{N}$, with ${\boldsymbol{s}}\in\mathbb{S}^{d-1}$. We focus on space dimensions $d=2,3$. Doing so leads to a countably-parametric description of the ensemble of admissible shapes, with parameters being the expansion coefficients.
Specifically, we adopt here the Karhunen-Loève-type shape expansion
$r(\boldsymbol{Y}(\omega);{\boldsymbol{s}})=\frac{1}{k}\,\sum\limits_{j=1}^{\infty}\beta_{j}\,Y_{j}(\omega)\,r_{j}({\boldsymbol{s}})\;,\quad{\boldsymbol{s}}\in\mathbb{S}^{d-1}\;,$ (7.1)
with $k$-independent weights $\beta_{j}>0$ and globally supported, real-valued expansion functions $r_{j}\in C^{\infty}(\mathbb{S}^{d-1})$. We assume that these are normalized such that $\left\|{r_{j}}\right\|_{C^{0}(\mathbb{S}^{d-1})}=1$, for all $j\in\mathbb{N}$. We confine the discussion to the following particular choices:
  • For $d=2$ we opt for
    $r_{j}(s):=\begin{cases}\sin(\tfrac{j}{2}\,s)&\text{for even }j\;,\\ \cos(\tfrac{j-1}{2}\,s)&\text{otherwise }\;,\\ \end{cases}\quad 0\leq s<2\pi\;.$ (7.2a)
    Here ${\left\|{r_{j}}\right\|_{C^{p,1}(\mathbb{S}^{1})}\sim j^{p+1}}$ as $j\to\infty$.
  • For $d=3$ the $r_{j}$ are rescaled real spherical harmonics:
    $r_{j}:=Y_{\ell,m},\;\;j=\ell^{2}+\ell+m+1,\;\;-\ell\leq m\leq\ell,\;\;\ell\in\mathbb{N}_{0}.$ (7.2b)
    Here $\left\|{r_{j}}\right\|_{C^{p,1}(\mathbb{S}^{2})}\sim j^{(p+1)/2}$ as $j\to\infty$.
We suppose that in (7.1),
  • $Y_{j}=Y_{j}(\omega)$ i.i.d uniformly in $[-1,1]$: $Y_{j}\sim\mathcal{U}([-1,1])$, and
  • the deterministic weight sequence $(\beta_{j})_{j\in\mathbb{N}}\in\ell_{1}(\mathbb{N})$ with $\left\|{(\beta_{j})}\right\|_{\ell_{1}(\mathbb{N})}\leq 1\leq k/3$.
These assumptions ensure $r\in{\cal R}$, cf. (2.1).
Remark 7.1.
The scaling with ${k^{-1}}$ and the assumption $Y_{j}\sim\mathcal{U}([-1,1])$ in (7.1) limits the shape variations in (7.1) to size $O(k^{-1})$ for $k\to\infty$, that is, to a size proportional to the wavelength when $\beta_{j}$ is bounded independent of $k$, which is a stronger requirement than the scaling $\left\|{(\beta_{j})}\right\|_{\ell_{1}(\mathbb{N})}\leq\frac{k}{3}$ stipulated above. In Section 1.4 we found this to be necessary for the validity of polynomial surrogate modeling, which we have in mind throughout this work.
Remark 7.2 (Shape parametrization. Radial expansion functions).
It is natural to assume that the shape variations of $D$ are invariant under rotations, from which we conclude $\operatorname{Cov}r({\boldsymbol{s}},{\boldsymbol{s}}^{\prime})=g({\boldsymbol{s}}\cdot{\boldsymbol{s}}^{\prime})$ for some covariance function $g:[{-1,1}]\to\mathbb{R}$. Then Karhunen-Loève expansion of $r=r(\omega)$ will yield exactly the radial spherical harmonic expansion functions $r_{j}$ of (7.2).
Remark 7.3.
All that follows can be adapted also to uncertainty parametrization with locally supported function systems, such as splines or wavelet functions; see §9 for further discussion.
For UQ one replaces the random variables $Y_{j}$ with deterministic parameters, and places a product probability measure on the set of parameters. In light of (7.1) the radial displacement function becomes a deterministic, affine-linear function on the parameter set
${\cal P}:=[{-1,1}]^{\mathbb{N}}=\left\{{\mathbf{y}}=\left(y_{j}\right)_{j\in\mathbb{N}}:\;-1\leq y_{j}\leq 1,\,\text{ for all }j\in\mathbb{N}\right\}\subset\ell^{\infty}(\mathbb{N})\;,$ (7.3)
via
$\boxed{r({\mathbf{y}};{\boldsymbol{s}}):=\frac{1}{k}\sum\limits_{j=1}^{\infty}\beta_{j}\,y_{j}\,r_{j}({\boldsymbol{s}})\;,\quad{\boldsymbol{s}}\in\mathbb{S}^{d-1}\;,\quad{\mathbf{y}}=(y_{j})_{j\in\mathbb{N}}\in{\cal P}}\;.$ (7.4)
[…]
$\breve{u}:{\cal P}\to H^{1}(B_{2})\quad,\quad{\mathbf{y}}\mapsto\breve{u}({\mathbf{y}};\cdot):=\widehat{u}(r({\mathbf{y}});\cdot)\;.$ (7.5)
Since $r\mapsto\widehat{u}(r;\cdot)$ could be extended into the complex domain, we can extend all ${\mathbf{y}}\in{\cal P}$-dependent functions to complex-valued parameter sequences ${\mathbf{z}}$ contained in (a suitable superset of) $[{-1,1}]^{\mathbb{N}}+\rm i\mathbb{R}^{\mathbb{N}}\subset\ell^{\infty}(\mathbb{N},\mathbb{C})$. Thanks to the affine dependence (7.4) of $r({\mathbf{y}};\cdot)$ on ${\mathbf{y}}$, the domain of analyticity of the complex-parametric solution manifold
${\mathbf{z}}\mapsto{\breve{u}({{\mathbf{z}}};\cdot):=\widehat{u}(\mathfrak{r}({{\mathbf{z}}};\cdot);\cdot)}$ (7.6)
can immediately be read off Corollary 4.2. Here, we wrote $\mathfrak{r}({\mathbf{z}})$, ${{\mathbf{z}}\in\ell^{1}(\mathbb{N},\mathbb{C})}$, for functions defined by (7.4) with $z_{j}\in\mathbb{C}$ in place of the real-valued parameters $y_{j}\in[-1,1]$.
Corollary 7.4 (Domain of analyticity of ${\mathbf{z}}\mapsto\breve{u}({{\mathbf{z}}};\cdot)$).
The parametric mapping
${{\mathbf{z}}}\in\ell^{\infty}(\mathbb{N},\mathbb{C})\mapsto{\breve{u}({{\mathbf{z}}};\cdot):=\widehat{u}(\mathfrak{r}({\mathbf{z}});\cdot)}\in H^{1}(B_{2})\;,$ (7.7)
defined by combining (4.2) and (7.4) with (7.6), is holomorphic on
[…]
with the $k$-independent constants ${C_{\operatorname{Re}}},C_{\operatorname{Im}}>0$ as in Corollary 4.2.
A sufficient condition for the “complexified” radial displacement function $\mathfrak{r}({\mathbf{z}})$ from (7.4) to belong to class ${\mathcal{A}}_{p}$ as defined in (5.10), that is, a sufficient condition for the uniform wavenumber-explicit convergence estimates for PML-based $h$-FEM in Theorem 5.8 and for the approximate far-field pattern from Corollary 6.4, can be expressed in terms of complex-valued parameter sequences ${\mathbf{z}}$ in the set ${\cal H}_{p}$ for $p\in\mathbb{N}$, where
${\cal H}_{p}={\cal H}_{p}(C_{\operatorname{Re}},C_{\operatorname{Im}},k):=\left\{{\mathbf{z}}\in{\cal H}(C_{\operatorname{Re}},C_{\operatorname{Im}},k):\sum\limits_{j=1}^{\infty}|\beta_{j}||\operatorname{Re}z_{j}|\left\|{r_{j}}\right\|_{C^{p,1}(\mathbb{S}^{d-1})}\leq C_{\operatorname{Re}}k\right\}$ (7.9)
for $p\in\mathbb{N}$.
From Corollary 6.4, (6.9), and the triangle inequality, we conclude $k$-explicit, and $h$-uniform stability of the complex-parametric, approximate far-field.
Corollary 7.5 (Domain of analyticity of ${\mathbf{z}}\mapsto{\breve{u}^{\rm scat}_{\infty,\rm PML}({\mathbf{z}};\cdot)_{h}}$).
[…]
Thanks to Remark 5.5 we can also conclude $k$- and $h$-uniform parametric holomorphy of the PML/FEM-Galerkin solutions.
Corollary 7.6 (Uniform parametric holomorphy of PML/FEM-Galerkin solutions).
Under the assumptions of Theorem 5.8, for
  • ${\mathbf{z}}\mapsto\breve{u}^{\rm alt}({\mathbf{z}};\cdot)$, the outgoing solution of (6.6) transformed to the nominal domain, obtained from $\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)$ via the correspondence (7.5),
  • ${\mathbf{z}}\mapsto\breve{u}^{\rm alt}_{\rm PML}({\mathbf{z}};\cdot)$ corresponding similarly to the exact solution of the PML variational formulation (5.1) with the corresponding right-hand side, and
  • the parameterized PML/FEM-Galerkin approximation ${\mathbf{z}}\mapsto\breve{u}^{\rm alt}_{\rm PML}({\mathbf{z}};\cdot)_{h}$ derived from the solution of (5.9),
the following holds.
These parametric solution families are holomorphic as an $H^{1}_{k}(\Omega)$-valued map in a $h$ and $k$-independent set ${\cal H}_{p}(C_{\operatorname{Re}},C_{\operatorname{Im}},{k_{1}})\subset\ell^{\infty}(\mathbb{N},\mathbb{C})$ (with $k_{1}$ as in Theorem 5.4), and, uniformly on this set, satisfy the bound (6.10), with $\breve{u}^{\rm alt}_{\rm PML}({\mathbf{z}};\cdot)$ and $\breve{u}^{\rm alt}_{\rm PML}({\mathbf{z}};\cdot)_{h}$ in place of $\widehat{u}^{\rm alt}(\mathfrak{r};\cdot)$, correspondingly.
8 Computational Shape Uncertainty Quantification
The $k$ explicit, uniform w.r. to $h$ parametric holomorphy of the PML-Galerkin solution is the basis for the error analysis of efficient, deterministic computation of quantities of interest (QoIs). In a UQ-context these are, for example, the expectation $\mathbb{E}(\breve{u}^{\rm scat}_{\infty})$ and, possibly, higher order spatial correlation functions of the corresponding random field $\omega\mapsto\breve{u}^{\rm scat}_{\infty}(\omega):=\breve{u}^{\rm scat}_{\infty}({\mathbf{Y}}(\omega);\cdot)$. Such QoIs will be deterministic, smooth functions on $\mathbb{S}^{d-1}$. Since all random variables $Y_{j}(\omega)$ are uniformly distributed in $[-1,1]$, we find with the parametric, deterministic shape representation $r({\mathbf{y}};\cdot)$ from (7.4) the following expressions for the mean of the far-field pattern corresponding to $\breve{u}$ in (7.5), i.e.
$\mathbb{E}(\breve{u}^{\rm scat}_{\infty})({\boldsymbol{q}})$ $=\int\nolimits_{{\cal P}}\breve{u}^{\rm scat}_{\infty}({{\mathbf{y}}};{\boldsymbol{q}})\,\mathrm{d}\mu({{\mathbf{y}}})\;,$ ${\boldsymbol{q}}\in\mathbb{S}^{d-1}\;,$ (8.1a)
for all far-field directions ${\boldsymbol{q}}=\nicefrac{{{\boldsymbol{x}}}}{{|{\boldsymbol{x}}|}}\in\mathbb{S}^{d-1}$. Here, the measure $\mu$ denotes the countable product of the uniform probability measure on $[-1,1]$. We add that in actual computations in (8.1) the far-field pattern $\breve{u}^{\rm scat}_{\infty}$ has to be replaced with the computable approximation $\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;\cdot)_{h}$ in (6.11).
The formula (8.1a) is a so-called “ensemble average” over all admissible scatterer shapes, and involve “infinite-dimensional integrals”, whose efficient approximation by (possibly higher-order and deterministic) quadrature formulas is addressed next.
8.1 High-Dimensional Smolyak / Sparse-Grid Quadrature
The QoI (8.1a) being a countably-parametric, deterministic integral, we discuss two classes of numerical integration: first, so-called Smolyak-Quadrature (see [66] and references there) and second, so-called Higher-Order QMC Quadrature, in particular so-called interlaced polynomial lattice rules (“IPL-QMC integration”, see [26]). In these references, convergence rates of both numerical integration methods have been proved to be independent of the dimension of the domain of integration, under suitable quantified holomorphy of the integrands. In view of the preceding discussion on quantified, wavenumber-explicit holomorphy of the parametric solutions $\{\breve{{\mathbf{y}}}\mapsto\hat{u}(\breve{{\mathbf{y}}}):\breve{{\mathbf{y}}}\in{\cal P}\}$, we connect the parametric holomorphy analysis to [26, 66]. We start by recalling a suitable concept of quantified parametric holomorphy, which implies dimension-independent convergence rates of (i) Smolyak-quadratures [65, 66], (ii) sparse-grid interpolants [14, 65] and (iii) higher-order Quasi-Monte Carlo quadratures [26].
8.1.1 $({\boldsymbol{b}},\bar{p},\varepsilon)$-Holomorphy
We quantify the domain of parametric holomorphy of the (analytic continuation of the) affine-parametric function $\{\breve{{\mathbf{y}}}\in{\cal P}\mapsto u_{\breve{{\mathbf{y}}}}:\breve{{\mathbf{y}}}\in{\cal P}\}$. We still work with the $k^{-1}$-scaling and the affine-parametric parametrizations $\breve{{\mathbf{y}}}\mapsto\mathfrak{r}(\breve{{\mathbf{y}}})$ from (7.4) and with the globally supported radial expansion functions $r_{j}$ from (7.2).
We present quantified parametric holomorphy for a generic “target” Banach space $X$. Observe that the QoI’s in (8.1a) take values in the separable, complex Banach space $X\in\{C^{0}(\mathbb{S}^{d-1};\mathbb{C}),L^{2}(\mathbb{S}^{d-1};\mathbb{C})\}$.
Definition 8.1 ($({\boldsymbol{b}},\bar{p},\varepsilon)$-Holomorphy).
Let $X$ be a complex Banach space with norm $\|\circ\|_{X}$. For $\varepsilon>0$, a sequence ${\boldsymbol{b}}=(b_{j})_{j\geq 1}\in(0,\infty)^{\mathbb{N}}$ and some $p\in(0,1)$, the parametric map
${\cal P}\ni\breve{{\mathbf{y}}}\mapsto u_{\breve{{\mathbf{y}}}}\in X$
is called $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphic if the following conditions hold.
  • (i)
    The map ${\cal P}\ni\breve{{\mathbf{y}}}\mapsto u_{\breve{{\mathbf{y}}}}$ is uniformly bounded, i.e. there exists a bound $M_{0}>0$ such that
    $\sup_{\breve{{\mathbf{y}}}\in{\cal P}}\|u_{\breve{{\mathbf{y}}}}\|_{X}\leq M_{0}\;,$
  • (ii)
[…]
    $\sum_{j\geq 1}(\rho_{j}-1)b_{j}\leq\varepsilon\;,$ (8.2)
    the map ${\cal P}\ni\breve{{\mathbf{y}}}\mapsto u_{\breve{{\mathbf{y}}}}$ admits a complex extension $\breve{{\mathbf{z}}}\mapsto u_{\breve{{\mathbf{z}}}}$ that is continuous w.r. to $\breve{{\mathbf{z}}}$ and holomorphic w.r. to each variable $\breve{z}_{j}$ on a cartesian product set of the form
[…]
    Here, for $j\geq 1$, ${\cal O}_{\rho_{j}}\subset\mathbb{C}$ is some open set with $[-1,1]\subset{\mathcal{D}}_{\rho_{j}}\subset{\cal O}_{\rho_{j}}$ with strict inclusions where, for $\rho>1$, ${\mathcal{D}}_{\rho}:=\{z\in\mathbb{C}:|z|\leq\rho\}$ denotes the closed disc in $\mathbb{C}$ of radius $\rho>1$.
  • (iii)
[…]
The significance of $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy lies in the fact that holomorphic maps between (complex) Banach spaces become, upon adopting affine-parametric representation of their arguments w.r. to a suitable representation system (such as, e.g. $\{\psi_{j}\}_{j\geq 1}$ with $\psi_{j}\sim\beta_{j}r_{j}$ where $r_{j}$ is as in Remark 7.2) $({\boldsymbol{b}},\varepsilon)$-holomorphic maps in terms of the coefficient sequences in the representation of inputs (see [66, Lem. 3.3]). The summability exponent $\bar{p}\in(0,1)$ of the sequence ${\boldsymbol{b}}$ determines the convergence rate of suitable Smolyak quadratures in (8.8) below.
8.1.2 Estimating the Smolyak Quadrature Error
We recall the definition of a sparse-grid Smolyak quadrature: given a sequence $(\chi_{n;j})_{j=0}^{n}$, $n\in\mathbb{N}_{0}$ of $n$-tuples of pairwise distinct points in $[-1,1]$, Smolyak quadratures are built on corresponding univariate interpolatory quadrature rules $(Q_{n})_{n\geq 0}$ with nodes $\chi_{n;0},...,\chi_{n;n}\subset[-1,1]$ and corresponding weights $w_{n;j}>0$ w.r. to the uniform (probability) measure $\frac{1}{2}\lambda^{1}$, i.e.
$Q_{n}f=\sum_{j=0}^{n}w_{n;j}f(\chi_{n;j})\;,\quad w_{n;j}:=\frac{1}{2}\int_{-1}^{1}\prod_{i=0,i\neq j}^{n}\frac{y-\chi_{n;j}}{\chi_{n;i}-\chi_{n;j}}dy\;.$
By construction, $Q_{n}$ is exact for univariate polynomials of degree $n$. The weights $w_{n;j}$ can be negative, in general. For the error bound [66, Theorem 2.16] to hold, we assume stability of the univariate quadrature points in the sense that there exist constants $C,\vartheta>0$ such that
$\forall n\in\mathbb{N}_{0}:\quad\sup_{0\neq f\in C^{0}([-1,1])}\frac{|Q_{n}f|}{\|f\|_{C^{0}([-1,1])}}\leq C(n+1)^{\vartheta}$ (8.4)
We denote in the following the array of univariate sampling points
${\boldsymbol{\chi}}:=\left((\chi_{n;j})_{j=0}^{n}\right)_{n\in\mathbb{N}_{0}}\;.$ (8.5)
We adopt the convention that $\chi_{0;0}=0$, i.e. $Q_{0}$ corresponds to the midpoint rule, and also set $Q_{-1}:=0$. We refer to [17] and the references there for concrete constructions of such points.
Multivariate anisotropic quadratures are built from the univariate hierarchy $(Q_{n})_{n\geq 0}$ by tensorization. Let =${\mathcal{F}}=\{{\boldsymbol{\nu}}\in\mathbb{N}_{0}^{\mathbb{N}}:|{\boldsymbol{\nu}}|<\infty\}$ denote the set of finitely supported multiindices, and let $\Lambda\subset{\mathcal{F}}$ be a downward closed777 We recall (e.g. [15, Def. 1.1]) that an index set $\Lambda\subset{\mathcal{F}}$ is downward closed (“d.c.” for short) if ${\boldsymbol{\nu}}\in\Lambda$ and ${\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}$ implies ${\boldsymbol{\mu}}\in\Lambda$. Here, ${\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}$ means $\mu_{j}\leq\nu_{j}$ for all $j$. finite index set. Then for ${\boldsymbol{\nu}}\in{\mathcal{F}}$, we define the multivariate tensor product quadrature $Q_{\boldsymbol{\nu}}:=\bigotimes_{j\in\mathbb{N}}Q_{\nu_{j}}$. Then the Smolyak Quadrature for a d.c. set $\Lambda\subset{\mathcal{F}}$ is defined by
$Q_{\Lambda}:=\sum_{{\boldsymbol{\nu}}\in\Lambda}\bigotimes_{j\in\mathbb{N}}(Q_{\nu_{j}}-Q_{\nu_{j}-1})\;.$ (8.6)
In particular, $Q_{\Lambda}$ admits the representation
$Q_{\Lambda}:=\sum_{{\boldsymbol{\nu}}\in\Lambda}\iota_{\Lambda,{\boldsymbol{\nu}}}Q_{\boldsymbol{\nu}}\;,\quad\mbox{where}\quad\iota_{\Lambda,{\boldsymbol{\nu}}}:=\sum_{{\boldsymbol{e}}\in\{0,1\}^{\mathbb{N}}:{\boldsymbol{\nu}}+{\boldsymbol{e}}\in\Lambda}(-1)^{|{\boldsymbol{e}}|}\;.$
Based on this representation, one numerical evaluation of $Q_{\Lambda}$ requires accessing the integrand function in all points in the (finite) set
${\mathrm{pts}}(\Lambda,{\boldsymbol{\chi}}):=\left\{(\chi_{\nu_{j};\mu_{j}}:{\boldsymbol{\nu}}\in\Lambda,\iota_{\Lambda,{\boldsymbol{\nu}}}\neq 0,{\boldsymbol{\mu}}\leq{\boldsymbol{\nu}}\right\}\subset{\cal P}\;.$
We refer to [66, Section 2.2] for details.
The main result from [66, Section 2.5] on the convergence rate of $Q_{\Lambda}$ for suitable downward closed sets $\Lambda\subset{\mathcal{F}}$ of “active quadrature orders” ${\boldsymbol{\nu}}\in\Lambda$ is as follows.
Theorem 8.2.
Let $Z$ and $X$ be complex Banach spaces. Denote for $r>0$ with $B^{Z}_{r}=\{\varphi\in Z:\|\varphi\|_{Z}<r\}$ the open ball in $Z$ centered at the origin of radius $r$. Assume that we are given a holomorphic map $\mathfrak{u}:B_{r}^{Z}\to X$, a real constant $\delta>0$, a sequence $(\psi_{j})_{j\in\mathbb{N}}\subset Z$, $r>0$, and $\bar{p}\in(0,1)$. Fix $\delta>0$ arbitrarily small. Assume further that the following hold:
  • (i)
    $\sum_{j\geq 1}\|\psi_{j}\|_{Z}<r$ and the sequence ${\boldsymbol{b}}=(\|\psi_{j}\|_{Z})\in\ell_{\bar{p}}(\mathbb{N})\subset\ell_{1}(\mathbb{N})$,
  • (ii)
    $\mathfrak{u}:B^{Z}_{r}\to X$ is holomorphic and bounded by $M_{u}$,
  • (iii)
    the collection of univariate quadrature abscissae ${\boldsymbol{\chi}}$ satisfies (8.4).
Define the countably parametric map $U:{\cal P}\to X$ via the composition
$\breve{{\mathbf{y}}}\mapsto U(\breve{{\mathbf{y}}}):=\mathfrak{u}\left(\sum_{j\geq 1}\breve{y}_{j}\psi_{j}\right).$ (8.7)
Then, there holds
  • (i)
    [$k$-independent parametric holomorphy] The parametric maps $U$ in (8.7) are $({\boldsymbol{b}},p,\varepsilon)$ holomorphic in the sense of Definition 8.1, with some $\varepsilon>0$ independent of $k$, and
  • (ii)
    [$k$-uniform convergence rate of Smolyak quadrature error] There exists a constant $C>0$ (depending on $\delta>0$, but independent of $k$) such that for every $\epsilon>0$, there exists a finite, downward closed multiindex set $\Lambda_{\epsilon}\subset{\mathcal{F}}$ with $|\Lambda_{\epsilon}|\to\infty$ as $\epsilon\to 0$ such that the following error bound holds:
[…]
This is [66, Thm. 2.16 and Thm.  4.3, Eqn.(4.1)], with the $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy followwing from [66, Lemma 3.3]. Inspection of the proofs in [66] reveals, in particular, that the constant $C>0$ in the quadrature error bound in the statement of [66, Thm.  4.3, Eqn.(4.1)] scales linearly in the integrand modulus $M_{u}$, whence the bound (8.8).
Here, as in (8.1a), the measure $\mu$ in (8.8) is a countable product probability measure obtained from the univariate scaled Lebesgue measure, $\frac{1}{2}\lambda^{1}$. For nested collections of univariate integration points in the sense of [66, Def. 2.1] it holds that $|{\mathrm{pts}}(\Lambda,{\boldsymbol{\chi}})|=|\Lambda|$ ([66, Lemma 2.1]),
The constant $C>0$ in (8.8) depends on the sequence ${\boldsymbol{b}}$ in the statement of the $({\boldsymbol{b}},\bar{p},\varepsilon)$ holomorphy of the parametric integrand function in (8.7).
8.2 Combined Smolyak-PML/FEM error bounds
With the Smolyak-quadrature error bound Thm. 8.2 in hand, we proceed to estimate the error in the quadrature-FE approximation of the corresponding mean of the far-field.
The integral (8.1a), is approximated by a Smolyak type quadrature formula $Q_{\Lambda}[\cdot]$ as in (8.6). We write, for any farfield direction ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$, and with the approximate farfield $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};{\boldsymbol{q}})_{h}$ as in (6.11),
$\mathbb{E}[\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}})]-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\\ =\mathbb{E}[\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}})-\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h})]+\mathbb{E}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\\ =:I+II.$ (8.9)
We estimate the $L^{2}(\mathbb{S}^{d-1};\mathbb{C})$-norms of the terms $I$ and $II$ in (8.9) separately.
Term $I$ is the expected discretization error of the PML-Galerkin FE discretization. For fixed direction ${\boldsymbol{q}}\in\mathbb{S}^{d-1}$, estimated by the uniform (w.r.t. all admissible shapes $\breve{y}\in{\mathcal{P}}$) error bound in the computed farfield pattern $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}$ which was bound in Corollary 6.4, (6.12). Due to $\mu({\mathcal{P}})=1$,
$\sup_{{\boldsymbol{q}}\in\mathbb{S}^{d-1}}\left|\mathbb{E}(\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}})-\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h})\right|\leq\sup_{\mathfrak{r}\in{\mathcal{A}}_{p}}\left\|{\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};\cdot)-\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}}\right\|_{L^{\infty}(\mathbb{S}^{d-1})}\;.$ (8.10)
We further majorize term $I$ by the bound (6.12), using the assumptions of Theorem 5.8 and with $k_{1}$ as in Theorem 5.4, for all $h,k$, and $p$ satisfying (5.11), provided that $\mathfrak{r}({\mathbf{y}};\cdot)\in{\mathcal{A}}_{p}$ uniformly w.r.t. ${\mathbf{y}}\in{\mathcal{P}}$.
A sufficient condition for the regularity $\mathfrak{r}({\mathbf{z}};\cdot)\in{\mathcal{A}}_{p}$ of the parametric geometry representation (7.2a)-(7.4) is that that ${\mathbf{z}}\in{\cal H}_{p}$, see (7.9). With the radial spherical harmonics expansion functions $r_{j}$ as in Rem. 7.2, i.e. (7.2a) for $d=2$ and (7.2b) for $d=3$, it holds that
$\|r_{j}\|_{C^{p+1}(\mathbb{S}^{d-1})}\simeq\|r_{j}\|_{C^{p,1}(\mathbb{S}^{d-1})}\simeq j^{(p+1)/(d-1)}\;,\;\;j=1,2,...\;.$
A sufficient condition for $\mathfrak{r}\in{\mathcal{A}}_{p}$ is that the weight sequence $(\beta_{j})_{j\geq 1}\in(0,\infty)^{\mathbb{N}}$ in (7.4) admits the bound
$\exists C>0:\quad\beta_{j}\leq Cj^{-1-\epsilon-(p+1)/(d-1)}\;,\;\;j=1,2,3,...\;.$ (8.11)
For ${\mathbf{y}}\in{\mathcal{P}}$, and every $p\geq 1$, the geometry representation in (7.1), i.e. the sum
$\sum_{j\geq 1}\beta_{j}y_{j}r_{j}(s)$
then converges in $C^{p,1}(\mathbb{S}^{d-1})\sim C^{p+1}(\mathbb{S}^{d-1})$, uniformly w.r. to ${\mathbf{y}}\in{\mathcal{P}}$, and (8.10) can be majorized by (6.12).
Term $II$ is the quadrature error of the Smolyak quadrature $Q_{\Lambda}$ in (8.6) applied to the parametric integrand $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}$. We estimate it with Theorem 8.2, assuming (8.11) on the weight sequence $\boldsymbol{\beta}$. We introduce the notation $\bar{p}\in(0,1]$ for the summability exponent of the weight sequence $\boldsymbol{\beta}$, and keep $p\in\mathbb{N}$ for the smoothness class in e.g. (7.9).
With $p\geq 1$, i.e. ${\mathbf{z}}\in{\cal H}_{p}$ as defined in (7.9), $\bar{p}$-summability of the weights $(\beta_{j})_{j\geq 1}\in(0,\infty)^{\mathbb{N}}$ satisfying (8.11) to bound term $I$ will imply, thanks to Theorem 8.2, quadrature convergence rates immune to the Curse of Dimensionality for Smolyak (or “sparse grid”) quadrature, as described in Section 8.1.2 and as analyzed in [66]. It will also imply, via the parametric $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy shown in corresponding results for higher-order Quasi-Monte Carlo (“HoQMC” for short) deterministic integration, as shown in [26, Thm. 3.1 and Prop. 4.1]. We leverage these results, with $k$-explicit error bounds, by verifying $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphy (cp. Def. 8.1) of the parametric integrands ${\mathbf{z}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}({\mathbf{z}};\cdot)_{h}$.
To verify the assumptions in Theorem 8.2, we remind that we work in the shape-parametrization (7.4) with the (globally supported, in $\mathbb{S}^{d-1}$) radial spherical harmonics expansion functions $r_{j}$ as in Rem. 7.2, i.e. (7.2a) for $d=2$ and (7.2b) for $d=3$. We choose in Theorem 8.2, item (i),
$Z=C^{2}(\mathbb{S}^{d-1};\mathbb{R})\subset C^{1,1}(\mathbb{S}^{d-1};\mathbb{R})\quad\mbox{and}\quad X=C^{0}(\mathbb{S}^{d-1};\mathbb{C})\;,$ (8.12)
with corresponding norms. The $C^{1,1}$-regularity implied by the choice of $Z$ is required in Theorem 5.8.
Based on the correspondence (7.6), we furthermore set $\psi_{j}=\beta_{j}r_{j}$ and identify $\mathfrak{u}$ in Theorem 8.2, item (ii), with $\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}$ in (6.11) resulting in the parametric integrand $U(\breve{{\mathbf{y}}})$ in (8.7) being $\breve{u}^{\rm scat}_{\infty,\rm PML}(\breve{{\mathbf{y}}};\cdot)_{h}$.
Assumption (8.11) on the weight sequence $\boldsymbol{\beta}=(\beta_{j})_{j\geq 1}$ and $\psi_{j}=\beta_{j}r_{j}$ implies with (7.2a), (7.2b) for $d=2,3$ with $p=2$ that (using $\|r_{j}\|_{C^{1,1}(\mathbb{S}^{d-1})}\sim\|r_{j}\|_{C^{2}(\mathbb{S}^{d-1})}\sim j^{3/(d-1)}$ as $j\to\infty$; the $C^{1,1}$-regularity is required by Theorem 5.8)
$b_{j}:=\|\psi_{j}\|_{Z}\simeq j^{1/(d-1)}\beta_{j}\lesssim j^{-(1+\epsilon+p/(d-1))}.$ (8.13)
Then, $(\|\psi_{j}\|_{Z})_{j\geq 1}\in\ell_{\bar{p}}(\mathbb{N})$, if
$\left(1+\frac{p}{d-1}\right)^{-1}\leq\bar{p}<1\;,\quad p=1,2,...$ (8.14)
The following result complements Cor. 7.5.
Proposition 8.3.
Suppose that the assumptions of Theorem 5.8 hold and assume (8.11) for the weight sequence $\boldsymbol{\beta}=(\beta_{j})_{j\geq 1}$, and the affine shape parametrization (7.4) with the expansion coefficients $\psi_{j}=\beta_{j}r_{j}$ with (7.2a), (7.2b) for $d=2,3$.
Then, there holds that the parametric PML-Galerkin FE farfield approximations
$\{{\mathbf{y}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}({\mathbf{y}};{\boldsymbol{q}})_{h}:{\mathbf{y}}\in{\mathcal{P}}\}$
as defined in (6.11) via (7.6) are, with the sequence ${\boldsymbol{b}}=(b_{j})_{j\geq 1}$ as defined in (8.13), $({\boldsymbol{b}},\bar{p},\varepsilon)$ holomorphic, uniformly w.r.t. $h$, $k$, in the space $X=C^{0}(\mathbb{S}^{d-1};\mathbb{C})$.
Proof.
This follows from [66, Lem. 3.3] with the summability exponent $\bar{p}$ bounded as in (8.14). ∎
With Prop. 8.3, the Smolyak error bound (8.8) then implies that with $p\geq 1$ as in (8.11)
$\begin{array}[]{rcl}\displaystyle\left\|\mathbb{E}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\right\|_{L^{2}(\mathbb{S}^{d-1};\mathbb{C})}&\leq&CM_{\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;\cdot)_{h}}|{\mathrm{pts}}(\Lambda_{\epsilon},{\boldsymbol{\chi}})|^{-\frac{2}{\bar{p}}+1+\delta}\\ &\leq&CM_{\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;\cdot)_{h}}|{\mathrm{pts}}(\Lambda_{\epsilon},{\boldsymbol{\chi}})|^{-1-\frac{2p}{d-1}+\delta}\end{array}$
By Corollary 7.5 and (7.10), the modulus $M_{\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;\cdot)_{h}}$ admits the $k$-explicit bound
$M_{\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;\cdot)_{h}}\leq\sup_{{\mathbf{z}}\in{\cal H}}\left\|{\breve{u}^{\rm scat}_{\infty,\rm PML}({\mathbf{z}};\cdot)_{h}}\right\|_{L^{2}({\mathbb{S}^{d-1}})}\leq Ck\|\widehat{u}^{\mathrm{inc}}\|_{H^{1}_{k}(B_{2})}\;.$
Combining this with the bound (8.10) for term $I$, we arrive for scatterer-geometries of regularity $C^{p,1}(\mathbb{S}^{d-1})$ for some $p\geq 1$ at the error bound
$\left\|\mathbb{E}(\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}}))-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\right\|_{L^{2}(\mathbb{S}^{d-1};\mathbb{C})}\\ \leq C\sup_{\mathfrak{r}\in{\mathcal{A}}_{p}}\left\|{\widehat{u}^{\rm scat}_{\infty}(\mathfrak{r};\cdot)-\widehat{u}^{\rm scat}_{\infty,\rm PML}(\mathfrak{r};\cdot)_{h}}\right\|_{L^{\infty}(\mathbb{S}^{d-1})}+Ck\|\widehat{u}^{\mathrm{inc}}\|_{H^{1}_{k}(B_{2})}|{\mathrm{pts}}(\Lambda_{\epsilon},{\boldsymbol{\chi}})|^{-1-\frac{2p}{d-1}+\delta}$ (8.15)
with the first term of the bound in turn majorized by (6.12). Absorbing the ($k$-independent) constants there, we arrive at the $k$-explicit error bound for the FE-PML and Smolyak-Quadrature approximated expected farfield pattern
$\left\|\mathbb{E}(\breve{u}^{\rm scat}_{\infty}(\cdot;{\boldsymbol{q}}))-Q_{\Lambda}[\breve{u}^{\rm scat}_{\infty,\rm PML}(\cdot;{\boldsymbol{q}})_{h}]\right\|_{L^{2}(\mathbb{S}^{d-1};\mathbb{C})}\\ \lesssim\|\widehat{u}^{\mathrm{inc}}\|_{H^{1}_{k}(B_{2})}\left[\big{(}hk+(hk)^{p}k\big{)}(hk)^{p}k+e^{-C_{\rm PML,2}k\big{(}R_{\rm tr}-(1+\epsilon)R_{1}\big{)}}+k|{\mathrm{pts}}(\Lambda_{\epsilon},{\boldsymbol{\chi}})|^{-1-\frac{2p}{d-1}+\delta}\right]$ (8.16)
Here, the constant hidden in $\lesssim$ is bounded independent of $k$, but depends on the regularity $p\geq 1$ of the geometry, and on the space dimension $d=2,3$.
Remark 8.4 (Higher Order Quasi-Monte Carlo - PML Galerkin FE error bound).
In Prop. 8.3, we observed that the parametric integrands ${\mathbf{y}}\mapsto\breve{u}^{\rm scat}_{\infty,\rm PML}(\breve{{\mathbf{y}}};\cdot)_{h}$ are $({\boldsymbol{b}},\bar{p},\varepsilon)$-holomorphic, uniformly with respect to $h$, as vector-valued integrand functions taking values in the (separable) Hilbert-space $X=L^{2}(\mathbb{S}^{d-1};\mathbb{C})$. For such integrand functions, certain deterministic so-called Quasi-Monte Carlo integration methods have been shown to furnish likewise higher orders of convergence, without the CoD. For general introduction and comprehensive presentation of this class of high-dimensional quadrature methods, we refer to [25].
Specifically, based on Prop. 8.3, it was shown in [26, Thm. 3.1, Prop. 4.1], that deterministic quadrature rules with $N$ points can be constructed that approximate the integral (8.1a) with rate $N^{-1/\bar{p}}$. The construction can be effected in $O(N\log N)$ many operations, subject to prior truncation to a suitable finite number $s$ of integration variables.
Particular QMC quadratures covered by the weighted norm setting in [26] allow for computable a-posteriori error estimation of the QMC integration error [27].
9 Conclusion
We have established shape-holomorphy of the approximate far-field pattern based on domain transformation, finite-element Galerkin approximation, and PML truncation for a class of time-harmonic acoustic scattering transmission problems. We have done this in the special setting of star-shaped scatterers parameterized by their radial extension and for piecewise homogeneous isotropic media, for which the refractive index outside the scatterer is larger than that inside. In this situation we could accomplish the first analysis that is fully explicit in the wavenumber $k$. We could also derive a $k$-explicit estimate for the sparse grid (Smolyak) quadrature error in the case of an affine parameterization of the radial displacement function that determines the shape of the scatterer.
In order to keep the focus of the article, several issues and possible extensions have not been addressed:
  • In the present affine-parametric representation (7.1)-(7.4), a basis $\{r_{j}({\boldsymbol{s}})\}_{j\geq 1}$ of $L^{2}(\mathbb{S}^{d})$ with global support in $\mathbb{S}^{d}$ was employed. Locally supported basis functions $r_{j}$ like (spline) wavelets for representing the radial displacement function are equally possible and offer advantages in terms of parsimonious representation of local features. The quantified holomorphy and quadrature convergence analysis for such representation systems can be performed along the present lines. We refer to [35] for details and computational comparisons, albeit in a $k$-implicit setting.
  • The investigation of multi-level algorithms for high-dimensional quadrature [24, 65]. The mathematical justification of these algorithms in the present setting requires, however, uniform with respect to $h$ and $k$ holomorphic dependence of the parametric solution $\widehat{u}(\mathfrak{r};\cdot)$ in norms which are stronger than $H^{1}_{k}(B_{2})$ in (4.5). In Theorem 4.1, part (ii), we already provided such bounds in (4.6).
  • The most general case of large wavenumber-independent “$O(1)$” shape variations, cf. the discussion in Section 1.4. In this case we expect a $k$-dependent, pre-asymptotic phase of the convergence of standard Smolyak quadrature. In this phase we may observe a dimension-dependent rate up to parameter dimension $O(k^{d-1})$. Smolyak constructions based on wavenumber-dependent (Filon-type) univariate quadrature may provide a partial remedy here, as proposed in [63].
  • The construction of efficient sparse polynomial surrogate shape-to-far-field maps based on interpolation.
  • The extension of the approximation results of Section 8 to the variance of the far-field pattern. This will confront the difficulty that the arising integral will fail to feature an analytic integrand, since we have to integrate the squared modulus of a complex-valued function.
Acknowledgements
EAS was supported by EPSRC grant EP/R005591/1. This work was initiated when EAS was a Nachdiplom lecturer at ETH in Autumn 2021, supported by the Forschungsinstitut für Mathematik (FIM).